### v1

# URS: A Unified Neural Routing Solver for Cross-Problem Zero-Shot Generalization

Abstract
Multi-task neural routing solvers have emerged as a promising paradigm for their ability to solve multiple vehicle routing problems (VRPs) using a single model. However, existing neural solvers typically rely on predefined problem constraints or require per-problem fine-tuning, which substantially limits their zero-shot generalization ability to unseen VRP variants. To address this critical bottleneck, we propose URS, a unified neural routing solver capable of zero-shot generalization across a wide range of unseen VRPs using a single model without any fine-tuning. The key component of URS is the unified data representation (UDR), which replaces problem enumeration with data unification, thereby broadening the problem coverage and reducing reliance on domain expertise. In addition, we propose a Mixed Bias Module (MBM) to efficiently learn the geometric and relational biases inherent in various problems. On top of the proposed UDR, we further develop a parameter generator that adaptively adjusts the decoder and bias weights of MBM to enhance zero-shot generalization. Moreover, we propose an LLM-driven constraint satisfaction mechanism, which translates raw problem descriptions into executable stepwise masking functions to ensure solution feasibility. Extensive experiments demonstrate that URS can consistently produce high-quality solutions for more than 100 distinct VRP variants without any fine-tuning, which includes more than 90 unseen variants. To the best of our knowledge, URS is the first neural solver capable of handling over 100 VRP variants with a single model.
1 Introduction
The Vehicle Routing Problem (VRP) is an essential class of combinatorial optimization problems (COPs) with extensive applications in logistics and supply chain management (Tiwari & Sharma, 2023; Sar & Ghadimi, 2023). Solving VRPs efficiently is challenging due to their NP-hard nature. While exact solvers can produce optimal solutions, they often become computationally prohibitive for real-world scenarios. In recent decades, classical heuristic solvers (Helsgaun, 2017; Vidal, 2022) have achieved impressive performance within acceptable timeframes. Nevertheless, these methods require considerable domain expertise to design specialized rules for each routing problem. Given the growing diversity of VRP variants in real-world applications, manually crafting tailored rules for every case has become impractical.
In recent years, neural combinatorial optimization (NCO) methods have attracted substantial attention for their potential to reduce reliance on handcrafted rules while maintaining competitive solution quality (Bengio et al., 2021; Li et al., 2022a). This has led to the development of many high-performing neural routing solvers that can automatically learn implicit problem-specific patterns from data under different training paradigms such as supervised learning (SL) (Drakulic et al., 2023; Luo et al., 2023; 2025b), reinforcement learning (RL) (Bello et al., 2016; Kool et al., 2019; Zhou et al., 2024a), and emerging self-improved learning (SIL) (Luo et al., 2023; 2025a; Pirnay & Grimm, 2024). Although these solvers have demonstrated impressive performance on specific problems (Kwon et al., 2020; Kim et al., 2022; Zhou et al., 2025), they typically require architectural customization and per-problem retraining to accommodate the distinct constraints and features of different VRP variants. These limitations increase the overall training cost and hinder practical deployment on new problems.
Table 1: Comparison between our URS and existing neural solvers with multi-task learning.
Neural Solver Training Paradigm Routing Problems (#\#) Zero-shot Generalization Remarks
MTPOMO (Liu et al., 2024a) RL 16 $\checkmark$ Constraint Combination
MVMoE (Zhou et al., 2024b) RL 16 $\checkmark$ Constraint Combination
RouteFinder (Berto et al., 2024) RL 48 $\checkmark$ Constraint Combination
CaDA (Li et al., 2025a) RL 16 $\checkmark$ Constraint Combination
MTL-KD (Zheng et al., 2025a) RL+KD 16 $\checkmark$ Constraint Combination
TSP-FT (Lin et al., 2024) RL 3 $\times$ Adapter-based Fine-tuning
MTL-MAB (Wang et al., 2025a) RL 3 $\times$ Adapter-based Fine-tuning
GOAL (Drakulic et al., 2025) SL 9 $\times$ Adapter-based Fine-tuning
[…]
To address the challenge of cross-problem generalization, growing attention has been directed toward multi-task learning capable of handling diverse routing problems. As summarized in Table 1, existing efforts largely fall into two categories: 1) constraint combination-based multi-task learning and 2) adapter-based fine-tuning frameworks. In the first approach, VRP variants are treated as different combinations of constraint attributes, and a unified model is trained across these combinations to enable knowledge sharing for problems with seen constraints (Liu et al., 2024a; Zhou et al., 2024b; Li et al., 2025a; Berto et al., 2024; Zheng et al., 2025a). The second approach builds a shared model backbone for all problems and incorporates specific input/output adapters for each problem, thereby reducing the retraining costs (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a).
Despite the above advancements, existing methods still fall short in zero-shot generalization. The problem coverage of constraint combination-based approaches is inherently bounded by their manually specified constraint sets, while adapter-based approaches require fine-tuning and thus cannot perform zero-shot generalization. Moreover, both strategies fundamentally rely on the explicit enumeration of problems through a predefined set of problem tags. This practice is problematic because the constraint space of routing problems is open-ended and compositional, making any fixed tagging scheme fundamentally limited. More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which is precisely what NCO always aims to avoid. A detailed review of related work on single-task and multi-task learning is provided in Appendix A.
In this paper, we propose a powerful Unified Neural Routing Solver (URS) to significantly improve the cross-problem zero-shot generalization ability for NCO methods. Unlike existing approaches, we introduce a unified data representation (UDR) that replaces problem enumeration with data unification, thereby broadening problem coverage and reducing reliance on domain expertise. In addition, we propose a Mixed Bias Module (MBM) to better learn the geometric and relational biases inherent in various problems. Based on UDR, we further develop a parameter generator that adaptively adjusts decoder and bias weights of MBM to improve zero-shot generalization. Furthermore, to handle complex and unseen routing problems, we introduce a constraint satisfaction mechanism that leverages a large language model (LLM) to translate raw problem constraint descriptions into executable node masking functions for each construction step, thereby ensuring our URS generates feasible solutions for various problems.
We evaluate URS on more than 100 VRP variants, including more than 90 unseen variants. The results demonstrate that URS not only obtains competitive performance against specialized neural solvers on seen variants but also exhibits strong zero-shot generalization capabilities on a wide range of unseen problems. To the best of our knowledge, URS is the first neural solver capable of efficiently solving over 100 distinct VRP variants using a single model without any retraining or fine-tuning.
2 Preliminaries
In this section, we first introduce the definition of VRP, then provide an overview of recent constructive neural solvers for solution generation (Kool et al., 2019; Kwon et al., 2020).
2.1 Vehicle Routing Problems
Consider a VRP instance with an optional depot indexed by 0 and $n$ customers indexed by $\{1,2,\ldots,n\}$, such as in the Capacitated VRP (CVRP). The instance can be represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the node set $\mathcal{V}=\{v_{i}\}_{i=0}^{n}$ has a total size of $|\mathcal{V}|=1+n$ unless the variant has no depot (e.g., the Traveling Salesman Problem), and the edge set is $\mathcal{E}=\{e(v_{i},v_{j})\ |\ v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$. The travel cost between nodes is given by a distance matrix $\bm{D}=\{d_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$. In VRP, each node $v_{i}\in\mathcal{V}$ includes node coordinates $\{x_{i},y_{i}\}$ when available and problem-specific attributes (e.g., demands in CVRP). A feasible solution is a tour $\bm{\pi}=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ (i.e., a finite sequence), which is a permutation of $m$ nodes. Let $\Omega$ denote the feasible sequence set that satisfies the constraints, given an objective function $f(\bm{\pi}|\mathcal{G})$, we aim to search for a sequence $\bm{\pi}^{*}$ with the maximum objective, i.e., $\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\Omega}f(\pi|\mathcal{G})$. In most routing problems, the objective can be represented as the negative value of the total distance of a feasible sequence $\bm{\pi}$.
In this paper, we consider over 100 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD). Detailed definitions of these constraints are provided in Appendix B.
2.2 Constructive Neural Routing Solver
Most constructive NCO methods employ an encoder-decoder architecture (Luo et al., 2023; Kwon et al., 2020). Let learnable model parameters be denoted as $\bm{\theta}=\{\bm{\theta}_{enc},\bm{\theta}_{dec}\}$. Without loss of generality, we present the prevailing construction pipeline using AM (Kool et al., 2019). Given an instance $\mathcal{G}$, raw node features are first mapped by a linear projection to initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$. They are then processed by an encoder $\bm{\theta}_{enc}$ with $L$ attention layers to produce refined node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. At each decoding step $t$, the decoder $\bm{\theta}_{dec}$ sequentially selects a node to append to the current partial solution $\pi_{1:t-1}=(\pi_{1},\pi_{2},...,\pi_{t-1})$, where $\pi_{1},\pi_{t-1}\in\mathcal{V}$ are the first and last visited node, respectively. This sequential process continues until a complete solution is formed. To guarantee the feasibility of generated solutions, a masking function $\mathcal{M}_{t}$ sets the selection probabilities of the following nodes to -$\infty$ during the construction process: nodes that (1) are visited or (2) violate problem-specific constraints (e.g., capacity).
3 The Proposed Model: Unified Neural Routing Solver
In this section, we propose a powerful Unified Neural Routing Solver (URS), which can significantly improve cross-problem zero-shot generalization for NCO methods. The framework is illustrated in Figure 1, and its key components are elaborated below.
Figure 1: The overview of our URS with a 44‑node APDTSP instance as an example. (a) The solving pipeline of URS; (b) The proposed mixed bias module that incorporates multiple geometric priors, enabling it to efficiently learn the geometric and relational biases inherent in various problems.
3.1 Unified Data Representation
Existing multi-task neural solvers (Liu et al., 2024a; Zhou et al., 2024b; Drakulic et al., 2025; Wang et al., 2025a) fundamentally rely on the explicit problem enumeration through a predefined set of problem tags, which is challenging because the constraint space of routing problems is open-ended and compositional, making any fixed tagging scheme fundamentally limited. More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which is precisely what NCO always aims to avoid.
From a data perspective, despite the substantial diversity in constraints among routing problems, their instances share a common instance representation, e.g., all input instances under the current constraint combinations (Liu et al., 2024a; Zhou et al., 2024b) can be formulated as instances of either CVRP or CVRPTW. Leveraging this commonality, we propose a Unified Data Representation (UDR) to replace problem enumeration with data unification, thereby broadening problem coverage and reducing reliance on domain expertise. For an instance $\mathcal{G}$, we define the unified representation $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$, each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a position identifier, unified attribute set, and node-type indicator, respectively.
Position Identifier ($\bm{\rho}_{i}$)
We define a unified positional identifier for each node as $\bm{\rho}_{i}=\{\eta_{i},x_{i},y_{i}\}\in[0,1]^{3}$. The $\eta_{i}\sim\mathrm{Uniform}(0,1)\in\mathbb{R}^{1}$ is an additional sampled scalar, which is used to address asymmetric problems, following  Drakulic et al. (2023). For symmetric cases, we simply set $\eta_{i}=0$.
Unified Attribute Set ($\bm{\omega}_{i}$)
We define the unified attribute set as $\bm{\omega}_{i}=\{\delta_{i},\epsilon_{i},\mu_{i},\ e_{i},\ l_{i},\ s_{i}\}$, where they correspond to demand, prize, penalty, earliest arrival time, latest arrival time, and service time, respectively. The unification empowers the model to learn the relative importance of individual attributes in a shared embedding space. Furthermore, it supports effortless extension or ablation of attributes through simple zero-filling, which eliminates the need for changes to the model architecture.
Node-type Indicator ($\bm{\xi}_{i}$)
We provide an additional 5-dimensional multi-hot type tensor $\bm{\xi}_{i}\in\{0,1\}^{5}$ for each node, encoding: depot, pickup node, delivery node, node in sub-routes, node in open route. The nodes in sub-routes and open routes mean the solution $\pi$ can have sub-routes or be an open route. We introduce explicit structural roles via $\bm{\xi}_{i}$ that help the model generalize over diverse routing problems. Further ablation analysis concerning $\bm{\xi}_{i}$ can be found in Appendix K.1.
Multi-hot Problem Representation ($\bm{\lambda}_{i}$)
Through the UDR, each problem instance activates only a specific subset of the unified feature space. Thus, for any instance $\mathcal{G}_{i}$ drawn from different problems, a corresponding multi-hot problem representation $\bm{\lambda}{i}$ is obtained by indicating active (non-zero) feature slots with a value of $1$. This representation captures only the presence of features, in contrast to their raw values, and it subsequently provides conditional guidance to the position bias weight (see Equation 6) and the adaptive decoder parameters $\bm{\theta}_{dec}(\bm{\lambda})$ (see Equation 8 and Equation 9), thereby helping the model in distinguishing problem variants and refining its node selection process.
Instead of relying on a predefined problem taxonomy with a discrete set of problem tags, our proposed UDR decouples instance representation from constraint definitions and operates in a unified space characterized by both continuous and discrete features. It allows a single neural solver to address a much larger, open-ended space of VRP variants, as new problems can be seamlessly incorporated into this unified representation. Problem-specific constraints are delegated to the model-agnostic masking function $\mathcal{M}_{t}$. For example, for all capacity-constrained problems, only the remaining load is explicitly provided as input, while diverse additional constraints (e.g., time windows, duration limit, and backhaul) are enforced implicitly by masking infeasible candidate nodes during solution construction, rather than being fully enumerated in the decoder’s input. This process is further automated by an LLM-driven constraint satisfaction mechanism that generates problem-specific mask generators, which convert constraint handling from manual engineering to automated programming (see Section 4 for more details). Together, our proposed UDR enables a single model to handle various explored and even unexplored VRP variants while significantly reducing reliance on domain expertise. Detailed descriptions of $\bm{U}$ and $\bm{\lambda}$ are provided in Appendix C and Appendix D, respectively.
3.2 Model Architecture
As shown in Figure 1, we adopt AM (Kool et al., 2019) as our basic model. While UDR provides cross-problem consistency, different problems have diverse geometric priors. To efficiently learn the geometric and relational biases inherent in various problems and obtain better generalization performance, we propose two key enhancements: (1) Mixed Bias Modules (MBM) that integrate diverse geometric priors, and (2) adaptive parameter generators conditioned on the UDR. Their implementations are detailed below.
Embedding Layer
Given an instance $\mathcal{G}$, for each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, we project each component into $d$-dimensional embeddings through three separate linear transformations:
$\mathbf{h}_{i}^{(0)}=\bm{\rho}_{i}W_{\bm{\rho}}+\bm{\omega}_{i}W_{\bm{\omega}}+\bm{\xi}_{i}W_{\bm{\xi}},$ (1)
where $W_{\bm{\rho}}\in\mathbb{R}^{3\times d},\ W_{\bm{\omega}}\in\mathbb{R}^{6\times d},\ W_{\bm{\xi}}\in\mathbb{R}^{5\times d}$ are learnable matrices. Then we obtain a set of initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$ for all nodes in instance $\mathcal{G}$.
Encoder
The encoder $\bm{\theta}_{enc}$ with $L$ stacked attention layers transforms $H^{(0)}$ into advanced node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. The detailed process is provided in Appendix E.1. Unlike prior work, we replace the plain attention (Vaswani et al., 2017) with the proposed mixed bias module (MBM) that incorporates three structural bias matrices via three separate attention calculations: (1) an outgoing distance matrix $\bm{D}$; (2) an incoming distance matrix $\bm{D^{\mathrm{T}}}$; and (3) an optional relation matrix $\bm{R}=\{r_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$, where $r_{ij}=0$ if a predefined relation exists (e.g., pickup–delivery pairing in PDPs) and 1 otherwise, so that smaller values remain preferable. For the $\ell$-th layer, the MBM output $\hat{\mathbf{h}}_{i}^{(\ell)}$ can be expressed as:
$\bar{\mathbf{h}}_{i}^{(0)}=\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,N,\bm{D}_{i}\right)),$ (2)
$\bar{\mathbf{h}}_{i}^{(1)}=\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,N,\bm{D}_{i}^{\mathrm{T}}\right)),\\$ (3)
$\bar{\mathbf{h}}_{i}^{(2)}=\begin{cases}\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,\bm{R}_{i}\right))&\text{if}\ \bm{R}\neq\emptyset\\ \bm{0}&\text{otherwise},\end{cases},$ (4)
$\hat{\mathbf{h}}_{i}^{(\ell)}=\left[\hat{\mathbf{h}}_{i}^{(0)},\hat{\mathbf{h}}_{i}^{(1)},\hat{\mathbf{h}}_{i}^{(2)}\right]W^{O},$ (5)
where $\cdot$$\cdot$] denotes the horizontal concatenation operator, $W^{O}\in\mathbb{R}^{(3\times d)\times d}$ is a learnable matrices. For $\mathrm{Attention}(\cdot)$, we adopt the Adaptation Attention Free Module (AAFM) (Zhou et al., 2024a), which boosts the perception of diverse geometric patterns through an adaptation function $f(\alpha,N,d_{ij})$ with bias weight $\alpha$ (see Appendix F for implementation details). We generate $\alpha$ via a lightweight bias network BIAS($\bm{\lambda}$) whose input is the instance’s multi-hot representation $\bm{\lambda}$:
$\mathrm{BIAS}(\bm{\lambda})=\max\left(1,\ \left(\bm{\lambda}W_{1}+\mathbf{b}_{1}\right)W_{2}+\mathbf{b}_{2}\right),$ (6)
where $W_{1}\in\mathbb{R}^{|\bm{\lambda}|\times d},\ W_{2}\in\mathbb{R}^{d\times 1},\ \mathbf{b}_{1}\in\mathbb{R}^{d},\ \mathbf{b}_{2}\in\mathbb{R}^{1}$ are learnable parameters. We set the minimum value of the generated bias to $1$, resulting in faster model convergence. Due to the relation $r_{ij}$ is scale-independent, we remove scale $N$ in adaptation function for $\bm{R}$, and if $\bm{R}$ is absent, we substitute a zero vector for $\bar{\mathbf{h}}_{i}^{(2)}$. As shown in Figure 1(b), unlike MatNet (Kwon et al., 2021), which processes $\bm{D}$ and $\bm{D^{\mathrm{T}}}$ using two parallel attention layers, we mix them into one shared attention layer and let separate heads address different priors. This head‑level incorporation not only cuts architectural redundancy but also enhances model flexibility (e.g., integrating relation matrix), enabling URS can achieve impressive performance for various VRP variants.
Decoder
Figure 2: Comparison between our URS and existing adapter-based fine-tuning frameworks.
In URS, all parameters of $\bm{\theta}_{dec}(\bm{\lambda})$ are generated based on the multi-hot problem representation $\bm{\lambda}$. Following Kwon et al. (2020), given the first and last visited node embeddings $\mathbf{h}_{\pi_{1}}^{(L)}$ and $\mathbf{h}_{\pi_{t-1}}^{(L)}$, and an optional constraint state $C_{t}\in\mathbb{R}^{1}$ (e.g., remaining load), the context embedding $\mathbf{h}_{(C)}^{t}$ can be expressed as:
$\mathbf{h}_{(C)}^{t}=\left[\mathbf{h}_{\pi_{1}}^{(L)},\ \mathbf{h}_{\pi_{t-1}}^{(L)},\ C_{t}\right]W_{Q}(\bm{\lambda}),$ (7)
where $\cdot$$\cdot$] is the horizontal concatenation operator, $W_{Q}(\bm{\lambda})\in\mathbb{R}^{(2d+1)\times d}$ is a linear projection matrix conditioned on $\bm{\lambda}$. Missing $C_{t}$ is zero-padded (see Appendix E.3 for detailed settings). The new context embedding $\hat{\mathbf{h}}_{(C)}^{t}$ is obtained via $\mathrm{Attention}(\cdot)$ (Zhou et al., 2024a) on $\mathbf{h}_{(C)}^{t}$ and $H^{(L)}$:
$\hat{\mathbf{h}}_{(C)}^{t}=\mathrm{Attention}(\mathbf{h}_{(C)}^{t},H^{(L)}W_{K}(\bm{\lambda}),H^{(L)}W_{V}(\bm{\lambda}),\mathcal{M}_{t},f(\alpha,N,d_{t,i})),$ (8)
where $W_{K}(\bm{\lambda})\in\mathbb{R}^{d\times d}$, $W_{V}(\bm{\lambda})\in\mathbb{R}^{d\times d}$ are linear projection matrices conditioned on $\bm{\lambda}$. Finally, we compute the selection probability for each feasible node $p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})$ based on the improved compatibility with adaptation bias (Zhou et al., 2024a):
$p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})=\mathrm{softmax}(\zeta\cdot\text{tanh}(\frac{\hat{\mathbf{h}}_{(C)}^{t}(\mathbf{h}_{i}^{(L)})^{\mathrm{T}}}{\sqrt{d}}+f(\alpha,N,d_{t,i}))+\mathcal{M}_{t})$ (9)
where $\zeta$ is the clipping parameter. Inspired by Lin et al. (2022), we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ (see Appendix E.2 for implementation details). For $\alpha$ in Equation 8 and Equation 9, we generate them via $\mathrm{BIAS}(\bm{\lambda})$ in Equation 6. As shown in Figure 2, compared with adapter-based frameworks (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a), we provide a unified decoding entrance, which $\mathrm{WEIGHT}(\bm{\lambda})$ and $\mathrm{BIAS}(\bm{\lambda})$ maps a problem representation vector $\bm{\lambda}$ directly to all parameters of the decoder. This conditional weight generation removes problem‑specific adapters, enabling zero-shot generalization to unseen problems while maintaining solution quality.
4 LLM-driven Constraint Satisfaction Mechanism
Figure 3: The overview of LLM-driven Constraint Satisfaction Mechanism. (a) The generating pipeline of LLM-driven Constraint Satisfaction Mechanism for a general routing problem; (b) The illustration of mask generator utilization, handling the complex constraint.
While the proposed UDR enables URS to handle a wide range of routing problems, ensuring solution feasibility across diverse and potentially unseen constraints remains a major challenge. Traditional NCO approaches rely on manually designed masks $\mathcal{M}_{t}$ for each specific VRP variant Kool et al. (2019); Kwon et al. (2020); Luo et al. (2023), which becomes impractical when solving over 100 distinct variants as targeted by URS. To address this bottleneck, we propose an LLM-driven constraint satisfaction mechanism that automatically produces problem-specific mask generators. This mechanism transforms constraint handling from a manual engineering task into an automated program design task, thereby effectively supporting URS’s excellent zero-shot generalization capabilities.
As illustrated in Figure 3 (a), our constraint satisfaction mechanism operates through a systematic generate-and-validate pipeline. The process begins by prompting an LLM with both a problem-specific natural language constraint description and a standardized Python code template. The LLM then designs a set of candidate mask generators that implement the constraint logic. Each candidate is validated on a set of problem instances using a problem-specific checker to verify its reliability in producing feasible solutions. Invalid candidates are discarded, and the generation process iterates until a reliable mask generator is obtained.
Specifically, during validation, the candidate mask generator guides the stepwise solution construction. As illustrated in Figure 3 (b), at each step $t$, the mask generator takes the current state of the partial solution $\pi_{1:t-1}$ as input, which contains various information such as visited nodes, remaining load, etc. It then produces a binary mask $\mathcal{M}_{t}$ that identifies all permissible nodes. This mask is directly integrated into the URS’s selection probability computation in Equation 9 to ensure feasibility.
To efficiently steer this generate-and-test process, we formulate it as an optimization problem that maximizes the solution validity rate. The well-developed LLM-based automated algorithm design frameworks, such as EoH (Liu et al., 2024b), can be effectively applied to tackle this optimization problem. The optimization process terminates immediately once a reliable mask generator is found, which significantly reduces computational overhead compared to exhaustive search approaches. Furthermore, generated mask generators can be cached and reused for different instances sharing the same problem variant, thereby amortizing the generation cost. For more details about our LLM-driven constraint satisfaction mechanism, please see Appendix G.
5 Experiments
In this section, we comprehensively evaluate URS against classical and learning-based solvers across more than one hundred distinct VRP variants. For URS, we focus on: (1) performance on seen problems and (2) zero-shot generalization under unseen ones. To ensure comparability with prior neural routing studies, we adopt the 100-node setting as our primary evaluation scenario for URS. All experiments are conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB memory).
5.1 Experimental Setup
Problem Setting
We train URS on $11$ mixed VRP variants to ensure that each feature in UDR has been seen at least once, i.e., $\mathcal{P}=\{P_{i}\}_{i=1}^{11}$. Specifically, training problems and related data generations are as follows: (1) ATSP (Kwon et al., 2021);(2) TSP (Kool et al., 2019); (3) CVRP (Kool et al., 2019); (4) ACVRP (Kwon et al., 2021; Kool et al., 2019); (5) Orienteering Problem (OP) (Kool et al., 2019); (6) PCTSP (Kool et al., 2019); (7) PDTSP (Li et al., 2021); (8) CVRPTW (Zhou et al., 2024b); (9) OCVRP (Zhou et al., 2024b); (10) CVRPB (Zhou et al., 2024b), and (11) OCVRPTW (Zhou et al., 2024b). To evaluate the zero-shot generalization of URS, we construct 96 unseen variants to further highlight the broad applicability of URS. And each test dataset consists of 1000 instances. The oracle solver for each problem can be found in Appendix D.
Model & Training Setting
We use an embedding dimension $d$ of $128$ and a feed-forward layer dimension of $512$. We set the number of attention layers in the encoder to $12$. The clipping parameter $\zeta=50$ in Equation 9 for better training convergence (Jin et al., 2023). For training, URS is trained by the REINFORCE (Williams, 1992) algorithm, and the data is generated on the fly (see Appendix E.4 for more details). We train URS for $500$ epochs with $2\text{,}000$ batches per epoch, and the batch size is $128$. At each training step, we randomly select one problem $P_{i}$ from the problem set $\mathcal{P}$. We employ the AdamW optimizer (Loshchilov & Hutter, 2017) with an initial learning rate of $10^{-4}$, which is decayed by a factor of 0.1 at the 451451st epoch. Weight decay is set to $10^{-6}$. More details about the model training are provided in Appendix H.
Baseline
We compare URS with the following methods: (1)Classical Solvers: [^1] We run LKH3 (Helsgaun, 2017) with $10000$ trials and 1 run (Kool et al., 2019). For PyVRP and OR-Tools, we run them on a single CPU core with time limits of 20s (Li et al., 2025a). PyVRP (Wouda et al., 2024), LKH3 (Helsgaun, 2017), and OR-Tools (Perron & Furnon, 2023); (2)Neural Solvers: (i) specialized solvers that support multiple VRP variants: Sym-NCO (Kim et al., 2022), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), ICAM (Zhou et al., 2024a) and POMO (Kwon et al., 2020); (ii) general neural solvers capable of handling multiple variants [^2] Given the differences in problem variants and visible task training among the baseline models, we only compare methods with the same training setup. Thus, RF Berto et al. (2024) and CaDA (Li et al., 2025a) are not included in our experiments to ensure a fair comparison.: MTPOMO (Liu et al., 2024a), MVMoE (Zhou et al., 2024b), ReLD-MTL (Huang et al., 2025), GOAL-MTL (Drakulic et al., 2025); (iii) Solvers for less-explored variants: MatNet (Kwon et al., 2021), Heter-AM (Li et al., 2021).
Metrics and Inference
We report the average objective value (Obj.), optimality gap (Gap), and total inference time (Time) for each method. The optimality gap quantifies the discrepancy between the solutions generated by the corresponding methods and the near-optimal solutions, which are obtained using classical solvers. For most NCO baseline methods, we execute the source code provided by the authors using default settings. We compare the performance of URS to the relevant baselines for each problem as well as its single-task version (i.e., URS-STL). For symmetric instances, we report the best result with $\times 8$ instance augmentation, following Kwon et al. (2020). For asymmetric instances, we report the best result with $\times 128$ instance augmentation, following Kwon et al. (2021).
Table 2: Performance on 1K test instances of $11$ seen VRPs.
Method ATSP TSP OP PCTSP PDTSP -
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 2m 0.00% 6m 0.00% 1.5m 0.00% 1.2h 0.00% 9.8m -
Sym-NCO - 0.14% 5s 0.68% 15s 0.14% 16s - -
BQ 1.27% 1m 0.35% 7s 0.63% 7s - - -
LEHD - 0.59% 2s - - - -
POMO - 0.13% 5s - - - -
ICAM 3.05% 1m 0.15% 4s - - - -
MatNet 0.95% 6.5m - - - - -
Heter-AM - - - - 6.61% 5m -
GOAL-MTL 1.77% 1m - 1.20% 38s - - -
[…]
URS 2.26% 1.1m 0.57% 6s 0.45% 4s 1.06% 5s 4.98% 4s -
Method ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 9.1m 0.00% 19.6m 0.00% 20.8m 0.00% 5.3m 0.00% 20.8m
Sym-NCO - 1.46% 6s - - - -
BQ - 3.24% 8s - - - -
LEHD - 3.68% 2s - - - -
[…]
ICAM - 2.04% 4s - - - -
MVMoE - 1.65% 12s 4.90% 15s 1.28% 11s 3.14% 13s 3.85% 15s
MTPOMO - 1.85% 6s 5.31% 8s 1.67% 5s 3.46% 6s 4.41% 7s
[…]
GOAL-MTL - 4.22% 48s 4.66% 42s - - -
[…]
URS 3.06% 1.4m 1.81% 6s 6.13% 8s 1.46% 6s 3.24% 6s 5.07% 8s
Table 3: Zero-shot generalization on 1K test instances of $30$ unseen VRPs. Due to page limit, complete experimental results (including all 96 unseen variants) are deferred to Appendix I.
Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 13.95% 14s 17.36% 14s 13.38% 14s 2.29% 15s 10.46% 16s
MTPOMO 13.71% 7s 16.94% 8s 13.44% 8s 3.26% 9s 11.35% 10s
ReLD-MTL 13.57% 9s 14.83% 9s 12.96% 10s 2.27% 11s 10.80% 12s
[…]
Method MDCVRP MDCVRPL MDOCVRPB MDOCVRPBPTW MDOCVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 31.90% 31s 33.42% 44s 28.21% 26s 46.18% 40s 45.23% 40s
MTPOMO 23.75% 27s 24.32% 25s 25.06% 20s 34.19% 29s 33.37% 31s
ReLD-MTL 17.37% 36s 18.48% 28s 18.50% 21s 34.08% 33s 33.83% 36s
[…]
Method ACVRPB ACVRPL ACVRPTW AOPDCVRP ACVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPBPTW ACVRPBPLTW ACVRPBPTW AOCVRPBPLTW ACVRPBTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDCVRPL AMDCVRPBPLTW AMDOCVRPBPLTW PDCVRP OPDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPBPTW AMDCVRPBPTW AMDCVRP APDTSP APDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
5.2 Performance Evaluation
Performance on Seen VRPs
The experimental results on trained VRPs with scale $100$ are reported in Table 2. We observe that among all comparable neural solvers, URS-STL achieves the best performance on 8 out of the 11 routing problems, complemented by remarkably fast inference, highlighting the strength of the URS architecture when fully specialized. Meanwhile, mixing all problems into the training of URS incurs only a minor performance drop, yet still delivers competitive performance across all problems, demonstrating that potential connections among different problems are effectively captured by URS.
Generalization to Unseen VRPs
We evaluate our URS model on zero-shot generalization across 96 unseen VRP variants, including 44 well-studied VRP variants from the NCO community and 52 newly introduced variants that have not been explored before. The detailed results are presented in Table 3 and Appendix I. On widely-studied constraints (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions. Notably, URS merely selects the next node from a given feasible candidates without any additional domain knowledge, making its performance especially meaningful. Benefiting from our UDR, URS can directly address previously unexplored variants while maintaining solution quality, which compounds asymmetry and PD relations. Without any fine-tuning, URS still constructs high-quality solutions for more than 90 VRP variants that are not seen in training, highlighting its strong zero-shot generalization ability.
Results on Benchmark Dataset
We further evaluate the generalization ability of URS on benchmark instances from CVRPLIB Set-X (Uchoa et al., 2017) with scale $\in[500,1000]$, and we compare URS against several representative single-task and multi-task models. As demonstrated in Table 12 (Appendix J), URS maintains its position as the best overall performance across instances of varying scales. Notably, RL-based URS surpasses the SL-based specialized model LEHD (8.678% vs. 12.836%), underscoring its practical applicability in real-world scenarios.
6 Ablation Study
Table 4: Effects of different components of URS.
Method Avg.gap
w/o $\bm{\xi}$ 3.78%
w/o $C_{t}$ 2.97%
$\{\bm{D}\}$ 6.33%
$\{\bm{D},\bm{D^{\mathrm{T}}}\}$ 4.07%
$\{\bm{D},\bm{R}\}$ 7.34%
POMO-Dec. 3.39%
ReLD-Dec. 3.87%
[…]
In this section, we conduct an ablation study across 11 seen VRP variants to validate the effect of key components of URS, mainly including: (1) without node-type indicator $\bm{\xi}$ of UDR; (2) without problem state $C_{t}$; (3) effects of different priors in MBM ; (4) effects of alternative decoder architectures. As detailed in Table 4. URS maintains its competitive performance without $C_{t}$. And its effectiveness is further improved by the inclusion of a node-type indicator $\bm{\xi}$. For MBM, the results reveal that three priors $\{\bm{D},\bm{D^{\mathrm{T}}},\bm{R}\}$ are complementary, and their combined use allows MBM to effectively learn the geometric and relational biases inherent in various problems, resulting in superior overall performance. Furthermore, replacing the decoder with POMO or ReLD confirms the strong cross-problem robustness of URS, while the integration of the decoder in ICAM (Zhou et al., 2024a) leads to a further improvement. More detailed results and analyses can be found in Appendix K.
7 Conclusion, Limitation, and Future Work
In this work, we propose a novel RL-based Unified Routing Solver (URS) for cross-problem zero-shot generalization. The core of URS is a Unified Data Representation (UDR), which broadens problem coverage by substituting problem enumeration with data unification. Its generalization capability is further enhanced by a Mixed Bias Module (MBM) and an adaptive parameter generator. Moreover, an LLM-based constraint satisfaction mechanism guarantees solution feasibility and reduces dependency on domain expertise. Extensive experiments demonstrate that URS can produce high-quality solutions for more than 100 distinct VRP variants without any fine-tuning, more than 90 of which are unseen during training.
Limitation and Future Work
While URS achieves impressive results on a wide range of VRP variants, its current training scheme does not accommodate their inherent differences in complexity. In the future, we aim to explore more effective training schemes to further improve training efficiency.
Appendix A Related Work
A.1 Neural Routing Solvers with Single-Task Learning
The NCO methods have emerged as a promising paradigm for tackling various routing problems, achieving impressive empirical performance. Early seminal work based on Pointer Networks ignited interest in learning end-to-end construction policies for TSP and CVRP (Vinyals et al., 2015; Bello et al., 2016; Nazari et al., 2018). Subsequently, the Transformer-style heavy-encoder and light-decoder auto-regressive architecture has become the dominant paradigm (Kool et al., 2019), while POMO (Kwon et al., 2020) exploits permutation symmetry to enhance exploration diversity. Many advancements within this paradigm have pushed promising results on small instances (e.g., 100-node TSPs) (Xin et al., 2020; Kim et al., 2022; Xiao et al., 2024a; b; Sun et al., 2024). However, models trained on a fixed instance scale often degrade sharply when evaluated out of scale. To mitigate the poor generalization performance, recent studies introduce auxiliary local policies (Gao et al., 2024), adaptive perception across instance scales (Zhou et al., 2024a), or different search space reduction methods (Fang et al., 2024; Zhou et al., 2025; Chen et al., 2025). In parallel, ”heavy decoder” designs that dynamically recompute embeddings at each construction step (Drakulic et al., 2023; Luo et al., 2023) demonstrate strong generalization on large-scale instances. To improve the performance on large-scale routing instances, complementary directions are reducing solving difficulty via problem decomposition (Zheng et al., 2024b; Ye et al., 2024b; Li et al., 2025b) or adopting non-autoregressive generation augmented with additional searches (e.g., Monte Carlo tree search (MCTS) and 2-opt) (Fu et al., 2021; Sun & Yang, 2023; Qiu et al., 2022; Li et al., 2023; Ye et al., 2023). Moreover, heterogeneous vehicle routing (Li et al., 2022b), asymmetric distance settings (Kwon et al., 2021), multiple VRPs (Zheng et al., 2024a), and pickup–delivery problems (Li et al., 2021; Kong et al., 2024) have each motivated tailored neural designs, which have yielded competitive results. Despite these advancements, they still train a specialized model per problem type, limiting cross-problem transfer. We instead focus on a unified neural routing solver that aims to generalize across multiple routing problems without retraining from scratch for each specification.
A.2 Neural Routing Solvers with Multi-Task Learning
To address the challenge of cross-problem generalization, growing attention has shifted to multi-task learning capable of adapting to diverse routing problems. Existing efforts can broadly fall into two categories: (1) constraint combination-based multi-task learning and (2) adapter-based fine-tuning frameworks. The former line (e.g., MTPOMO (Liu et al., 2024a)) treats VRP variants as combinations of a predefined set of attributes and trains a shared backbone, achieving promising results on up to $16$ VRP variants. Follow-up work extends this paradigm by introducing mixture-of-experts (MoE) modules (Zhou et al., 2024b), heavy decoder designs plus knowledge distillation (Zheng et al., 2025a), and constraint-aware dual-attention (Li et al., 2025a) or updated Transformer structure (Berto et al., 2024) to improve model performance. While this enables knowledge sharing across problems, its coverage is inherently bounded by the hand-specified attribute set. The second family trains a shared backbone and performs adaptation to each variant via lightweight adapters. For instance, Lin et al. (2024) pretrain a TSP backbone and adapt to new variants through adapter-based fine-tuning. Drakulic et al. (2025) integrate edge features and adopt supervised multi-task pretraining for obtaining a powerful backbone before fine-tuning on novel tasks. Wang et al. (2025a) propose a multi-armed bandit strategy to realize more efficient multi-task training. Although these reduce retraining cost, they fail to achieve zero-shot generalization. In contrast, we focus on a constructive neural routing solver supported with zero-shot capability: once trained, it can directly generate high-quality solutions for previously unseen problems without any fine-tuning.
A.3 LLM-based Algorithm Generation in VRPs
Employing LLMs to solve VRPs is a new research direction. As an early attempt, Yang et al. (2023) utilizes the LLM as a direct optimizer, prompting an LLM to directly generate solutions for TSP instances. A more predominant stream leverages LLMs as automated algorithm designers to generate algorithms for solving VRPs. These methods often integrate LLMs with a search procedure. EoH (Liu et al., 2024b) combined an LLM with Evolutionary Computation to design heuristic components for a Guided Local Search (GLS) framework. ReEvo (Ye et al., 2024a) incorporated a reflection mechanism to refine heuristics for iterative solution construction. HSEvo (Dat et al., 2025) additionally introduced a diversity mechanism for designing heuristics solving TSP and OP. And AHD-MCTS (Zheng et al., 2025b) employed Monte Carlo Tree Search (MCTS) to guide the design of heuristics for solving TSP and CVRP. Despite their promise, these methods share common limitations rooted in their reliance on the LLM’s implicit, pre-trained knowledge as the core design logic. The process of discovering effective algorithms demands extensive search procedures and is computationally expensive. Furthermore, the heuristics generated by the LLM are susceptible to overfitting to the specific instances used during the design phase, limiting their generalization capabilities. In contrast, our URS overcomes these challenges by establishing a clear division of labor: the LLM is tasked with satisfying complex constraints, while our URS model is responsible for ensuring solution quality. This approach significantly reduces the reasoning burden on the LLM, mitigating the risk of overfitting and enabling URS to effectively solve a more extensive range of VRPs.
Appendix B Setups of VRP Variants
In this paper, we consider over 100 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD). The node coordinates are randomly sampled from the unit square when available (in symmetric geometric cases, i.e., $d_{ij}=d_{ji}$), following AM (Kool et al., 2019). Below, we provide a comprehensive description of the data generation process for each constraint.
Capacity (C)
The selection of the vehicle capacity $\mathcal{C}$ is critical, as prior work (Luo et al., 2025c) demonstrates its significant impact on the solution structure. For PDVRP variants (e.g., PDCVRP, OPDCVRP, APDCVRP, and AOPDCVRP), an excessively large $\mathcal{C}$ causes the solution to approximate that of a PDTSP instance. Conversely, an overly restrictive $\mathcal{C}$ promotes a simplistic, repetitive Pickup-Delivery sequence. To ensure structurally sound solutions, we set $\mathcal{C}$=20 for PDCVRP variants. For all other VRP variants, we adopt the convention from AM (Kool et al., 2019), setting $\mathcal{C}$=50.
Open Route (O)
The open route setting implies that vehicles are not required to return to the depot after completing their service tours. This does not alter the data generation, and it specifically affects the final cost evaluation by excluding the travel distance from the last customer back to the depot.
Backhaul (B)
The standard CVRP involves vehicles delivering goods from a depot to a set of customers with positive demands. The introduction of backhauls extends this formulation to the CVRP with Backhauls (CVRPB), where vehicles must also load goods from backhaul customers and return them to the depot, in addition to serving the standard linehaul (delivery) customers. Following the experimental setup of MTPOMO (Liu et al., 2024a), we generate customer demands by sampling from a discrete uniform distribution over the integers $\{1,\ldots,9\}$. Subsequently, 20% of these nodes are randomly designated as backhaul customers, and their corresponding demand values are negative, while the remainder are classified as linehaul customers. Notably, no precedence constraints are imposed between the servicing of linehaul and backhaul customers.To ensure the feasibility of the solutions constructed by our neural solver, we adhere to the MVMoE (Zhou et al., 2024b) framework’s policy that the first customer visited must be a linehaul node.
Backhaul and Priority (BP)
This constraint imposes a strict precedence requirement on the standard CVRPB, dictating that all linehaul customers must be serviced prior to any backhaul customers. Consequently, once a vehicle serves a backhaul customer, it is prohibited from visiting any subsequent linehaul nodes within the same route. If the first customer served is a backhaul node, the vehicle’s load is initialized to zero. This ensures the neural solver can construct a feasible solution.
Duration Limit (L)
Following MTPOMO (Liu et al., 2024a), we impose a duration limit (i.e., the maximum length of each vehicle route) of 3 in symmetric cases. Since the node coordinates are randomly sampled from the unit square, the maximum possible distance between any two nodes is $\sqrt{2}$. The setting is sufficient to ensure a vehicle can deliver goods to at least one customer and return to the depot, thus guaranteeing the feasibility of the solution. For the setting of asymmetric cases, we provide details in the description of the following asymmetric instances.
Time Windows (TW)
Consistent with the MVMoE (Zhou et al., 2024b), we set time windows of the depot to $e_{0},l_{0}]$=[0,3] and its service time to 0. The service time $s_{i}$ for each customer node is set to 0.2. For the time windows for each customer node, we generate them following the methodology outlined in MVMoE (Zhou et al., 2024b).
Multi-Depot (MD)
The generation of coordinates for the multiple depots aligns with that of other nodes, with the number of depots set to 3 (Berto et al., 2024). A vehicle is required to return to the same depot it departed from. To guarantee that each starting node is considered in this multi-depot variant, we set the number of trajectories for an instance to be the product of the number of depots and the number of customer nodes, maximizing the exploration of high-quality solutions.
Prize Collecting (PC)
Under the Prize Collecting condition, each node (excluding the depot) is assigned both a prize and a penalty. A prize is collected for each visited node, while a penalty is incurred for each unvisited node. The objective is to minimize the total travel distance plus the sum of penalties for all unvisited nodes. A key constraint is that a minimum amount of total prize must be collected before the vehicle is permitted to return to the depot and terminate its route. Consistent with the AM (Kool et al., 2019) framework, the prizes and penalties are generated as follows: $prize\sim\text{Uniform}\left(0,\frac{4}{n}\right)$, and $penalty\sim\text{Uniform}\left(0,3\cdot\frac{k_{n}}{n}\right)$, where $n=100$ and $k_{n}=4$ in our experiments.
Asymmetric (A)
We adopt the data generation method from MatNet (Kwon et al., 2021) to create an asymmetric distance matrix. Due to the small pairwise distances in the generated distance matrix, the parameters for duration limit and time windows set for symmetric cases are not applicable to the asymmetric problems. To ensure the constraints remain valid, we adjust the duration limit and depot end time to 0.6 and 1.0, respectively.
Pickup and Delivery (PD)
For PDP variants, we adhere to the methodology in  Li et al. (2021). For a problem of size $N$, node 0 is set to the depot, nodes $1$ to $N/2$ are designated as pickup nodes, while nodes from $N/2+1$ to $N$ are designated as delivery nodes. For each pickup node $v_{i}$, its corresponding delivery node is $v_{i+N/2}$. A strict precedence constraint is imposed in PDP cases, requiring each vehicle to visit a pickup node before its corresponding delivery node. For PDCVRP variants, the demand for each delivery node is randomly sampled from a discrete uniform distribution over $\{1,2,\ldots,9\}$. The demand for each pickup node is then set to the negative of its corresponding delivery node’s demand. To ensure the feasibility of the constraints, the vehicle capacity is set to $20$.
Appendix C Unified Data Representation
In this section, we provide a detailed description of our UDR $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$ (see Table 5). Each node $\mathbf{u}_{i}$ consists of three components, i.e., $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$. where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a position identifier, unified attribute set, and node-type indicator, respectively. To distinguish linehaul/delivery from backhaul/pickup nodes, we assign negative demand values to backhaul/pickup nodes (positive for linehaul/delivery), making demand the only signed component, following  (Liu et al., 2024a; Zhou et al., 2024b).
Table 5: Detailed description of each component in UDR.
Attribute Name Symbol Range Description Examples
Random Identifier $\eta$ $[0,1]$ Random scalar for asymmetric problems. ATSP
Node Coordinates $\{x,y\}$ $0,1]^{2}$ 2D coordinates of the node. TSP, CVRP
Demand $\delta$ [1,1][-1,1] Node demand (++ for delivery, - for pickup). CVRP, PDCVRP
Prize $\epsilon$ $[0,1]$ Prize collected for each node. OP, PCTSP
Penalty $\mu$ $[0,1]$ Penalty incurred for each node. PCTSP
Earliest Arrival Time $e$ $[0,1]$ Earliest permissible arrival time. CVRPTW
Latest Arrival Time $l$ $[0,1]$ Latest permissible arrival time (deadline). CVRPTW
Service Time $s$ $[0,1]$ Time required for service at the node. CVRPTW
Depot - $\{0,1\}$ Indicates if the node is a depot. CVRP
Pickup Node - $\{0,1\}$ Indicates if the node is a pickup location. PDTSP
Delivery Node - $\{0,1\}$ Indicates if the node is a delivery location. CVRP, PDTSP
Node in Sub-routes - $\{0,1\}$ The solution $\pi$ can have sub-routes. CVRP
Node in Open Route - $\{0,1\}$ Vehicles need not return to the depot. OCVRP
Appendix D Multi-hot Problem Representation
Leveraging the UDR, each problem instance activates only a subset of the unified feature space. Thus, for any instance $\mathcal{G}_{i}$ from different problems, we simply mark active (non-zero) feature slots with 1 to obtain a corresponding multi-hot problem representation $\bm{\lambda}_{i}$. Different from raw feature values, it encodes presence only and provides conditional signals to the following position bias weight (see Equation 6) and adaptive decoder (see Equation 8 and Equation 9), helping the model distinguish variants and refine node selection. Here, we provide detailed representations for each problem in our experiments, and each problem is a 13-dimensional multi-hot vector.
As shown in Table 6, we train URS on eleven classic and varied routing problems to ensure that each feature in the UDR has been seen at least once, i.e., $\mathcal{P}=\{P_{i}\}_{i=1}^{11}$. Then, we evaluate generalization on unseen problems on different combinations of ten constraint attributes (see Appendix B), and the detailed representations of each unseen problem can be found in Table 7 (documented VRP variants) and Table 8 (undocumented VRP variants).
Table 6: The representations for seen routing problems in our experiments. The total number of problems is $1$. In training, each feature in the UDR has been seen at least once. Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
Problem Oracle RI Coord. Demand Prize Penalty EAT LAT ST Depot Pickup Delivery Sub-routes Open Route
ATSP LKH-3 $\checkmark$
TSP LKH-3 $\checkmark$
OP Compass $\checkmark$ $\checkmark$ $\checkmark$
PCTSP ILS $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PDTSP LKH-3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRP LKH-3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 7: The problem representations of $44$ VRP variants that have been extensively explored in current literature (Liu et al., 2024a; Berto et al., 2024). Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
Problem Oracle RI Coord. Demand Prize Penalty EAT LAT ST Depot Pickup Delivery Sub-routes Open Route
CVRPL LKH-3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPB PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPB PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SPCTSP ILS $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 8: The problem representations of $52$ VRP variants that have not yet been explored in current literature. Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
Problem Oracle RI Coord. Demand Prize Penalty EAT LAT ST Depot Pickup Delivery Sub-routes Open Route
ACVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
APDTSP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OPDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
APDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOPDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Appendix E Model Architecture
E.1 Encoder
Similar to previous work (Kool et al., 2019; Kwon et al., 2020), we also use the encoder composed of $L$ stacked attention layers, where each attention layer consists of two sub-layers: an attention sub-layer and a Feed-Forward (FF) sub-layer, both of which use Instance Normalization (Ulyanov et al., 2016) and skip-connection (He et al., 2016).
Let ${H}^{(\ell-1)}=\{\mathbf{h}_{i}^{(\ell-1)}\}_{i=0}^{n}$ denote the input to the $\ell$-th attention layer for $\ell=1,\ldots,L$. The outputs for the $i$-th node are computed as follows:
$\widetilde{\mathbf{h}}_{i}^{(\ell)}=\mathrm{IN}^{(\ell)}\left(\mathbf{h}_{i}^{(\ell-1)}+\mathrm{MBM}^{(\ell)}\left(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},\bm{D},\bm{D}^{\mathrm{T}},\bm{R}\right)\right),$ (10)
$\mathbf{h}_{i}^{(\ell)}=\mathrm{IN}^{(\ell)}\left(\mathbf{\widetilde{h}}_{i}^{(\ell)}+\mathrm{FF}^{(\ell)}\left(\mathbf{\widetilde{h}}_{i}^{(\ell)}\right)\right),$ (11)
where $\mathrm{IN}(\cdot)$ denotes instance normalization (Ulyanov et al., 2016), $\mathrm{MBM}$ represents the adopted mixed bias module (see Equation 5 for details), and $\mathrm{FF}(\cdot)$ in Equation 11 corresponds to a fully connected neural network with ReLU activation.
After $L$ attention layers, the final node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$ encapsulate the advanced feature representations of each node.
Notably, to ensure dimensional consistency across bias channels, we normalize $\bm{D}$ and $\bm{D^{\mathrm{T}}}$ by dividing their maximum value before passing the MBM.
E.2 Decoder
In URS, all parameters $\bm{\theta}_{dec}(\bm{\lambda})$ are adaptively generated conditioned on the multi-hot problem representation $\bm{\lambda}$ (see Appendix D for detailed problem representations). $\bm{\theta}_{dec}(\bm{\lambda})$ includes two MLP networks, which are $\mathrm{WEIGHT}(\bm{\lambda})$ and $\mathrm{BIAS}(\bm{\lambda})$. $\mathrm{WEIGHT}(\bm{\lambda})$ is used to generate linear projection matrices $W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})]$. $\mathrm{BIAS}(\bm{\lambda})$ in Equation 6 is used to generate bias weights for $\mathrm{Attention}(\cdot)$(see Equation 8) and compatibility (see Equation 9).
For $\mathrm{WEIGHT}(\bm{\lambda})$, we use a multi-layer MLP hypernetwork conditioned on a 13-dimensional problem representation $\bm{\lambda}$ to generate problem-specific adaptive parameters. Note that we factor $W_{Q}(\bm{\lambda})\in\mathbb{R}^{(2d+1)\times d}$ into three vertically concatenated blocks $W_{\text{first}}(\bm{\lambda})\in\mathbb{R}^{d\times d},W_{\text{last}}(\bm{\lambda})\in\mathbb{R}^{d\times d}$ and $W_{C}(\bm{\lambda})\in\mathbb{R}^{1\times d}$, applied respectively to the first visited node $\mathbf{h}_{\pi_{1}}^{(L)}$, the last visited node $\mathbf{h}_{\pi_{t-1}}^{(L)}$, and a scalar state feature $C_{t}\in\mathbb{R}^{1}$ (see Appendix E.3 for more details). Thus $\mathrm{WEIGHT}(\bm{\lambda})$ outputs the linear projection matrices $\{W_{\text{first}}(\bm{\lambda}),W_{\text{last}}(\bm{\lambda}),W_{C}(\bm{\lambda}),W_{K}(\bm{\lambda}),W_{V}(\bm{\lambda})\}$.
Furthermore, we adopt the parameter compression technique of Ha et al. (2016) to control model size. Specifically, we first generate an advanced embedding matrix $\bm{H}_{\text{dec}}(\bm{\lambda})=[\bm{h}_{\text{dec}}^{i}(\bm{\lambda})|i=1,2,\ldots,5]\in\mathbb{R}^{5\times|\bm{\lambda}|}$ whose five row embeddings each parameterize one target projection matrix. It can be computed via three consecutive linear projections:
$\bm{H}_{\text{dec}}(\bm{\lambda})=((\bm{\lambda}W_{1}+\mathbf{b}_{1})W_{2}+\mathbf{b}_{2})W_{3}+b_{3},$ (12)
where $W_{1}\in\mathbb{R}^{|\lambda|\times d_{h}},\ W_{2}\in\mathbb{R}^{d_{h}\times d_{h}},\ W_{3}\in\mathbb{R}^{d_{h}\times(5\times|\bm{\lambda}|)},\ \mathbf{b}_{1}\in\mathbb{R}^{d_{h}},\ \mathbf{b}_{2}\in\mathbb{R}^{d_{h}},\mathbf{b}_{2}\in\mathbb{R}^{(5\times|\bm{\lambda}|)}$ are learnable parameters, $d_{h}$ is $256$ in this work, following Lin et al. (2022). Next, we use five different linear projections to map the new representation $\bm{h}_{\text{dec}}^{i}(\bm{\lambda})$ of each row to the corresponding decoder parameters. The computation can be expressed as:
$W_{\text{first}}(\bm{\lambda})=\bm{h}_{\text{dec}}^{1}(\bm{\lambda})W^{1}_{\textbf{hyper}},\ W_{\text{last}}(\bm{\lambda})=\bm{h}_{\text{dec}}^{2}(\bm{\lambda})W^{2}_{\textbf{hyper}},\ W_{C}(\bm{\lambda})=\bm{h}_{\text{dec}}^{3}(\bm{\lambda})W^{3}_{\textbf{hyper}},$ (13)
$W_{K}(\bm{\lambda})=\bm{h}_{\text{dec}}^{4}(\bm{\lambda})W^{4}_{\textbf{hyper}},\ W_{V}(\bm{\lambda})=\bm{h}_{\text{dec}}^{5}(\bm{\lambda})W^{5}_{\textbf{hyper}},$ (14)
where $W^{1}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{2}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{3}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(1\times d)}$, $W^{4}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$, $W^{5}_{\textbf{hyper}}\in\mathbb{R}^{|\lambda|\times(d\times d)}$ are learnable parameters. In this way, we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate used linear projection matrices $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ conditioned on the multi-hot problem representation $\bm{\lambda}$, enabling the adaptive construction of high-quality solutions for different problems.
E.3 Problem State $C_{t}$
In this section, we detail the problems that require additional state signals. If a problem is not mentioned in this section, it means we will not apply any extra state to it, i.e., $C_{t}=0$.
For all problems with the capacity constraint, we explicitly impose only remaining load, while diverse additional constraints (e.g., time windows, duration limit, and backhaul) are enforced implicitly as $\mathcal{M}_{t}$ masks infeasible candidate nodes to guarantee valid solutions, instead of being fully enumerated in the input of the decoder.
For OP, we keep track of the remaining maximum length at time $t$, and $C_{t}$ represents the remaining routing length that can be moved, and we normalize it to $[0,1]$ by dividing by the maximum length (i.e., 4.0 in OP100). The setting is the same as AM (Kool et al., 2019).
For PCTSP and SPCTSP, $C_{t}$ is the remaining prize to collect, and we do not provide any information about the penalties, as this is irrelevant for the remaining decisions, following Kool et al. (2019).
For the problem state, we provide an ablation study in Appendix K.2.
E.4 Training
Following Kwon et al. (2020), we use $n$ trajectories with distinct starting nodes in training. URS is trained by the REINFORCE (Williams, 1992) algorithm with a shared baseline (Kwon et al., 2020):
$\nabla_{\theta}\mathcal{L}(\bm{\theta})=\mathbb{E}_{p(\bm{\pi}|\mathcal{G},{\bm{\theta}})}[(f(\bm{\pi}|\mathcal{G})-\frac{1}{n}\sum_{i=1}^{n}f(\bm{\pi}^{i}|\mathcal{G}))\nabla_{\bm{\theta}}\log p_{\bm{\theta}}\left(\bm{\pi}|\mathcal{G}\right)],$ (15)
$p_{\bm{\theta}}(\bm{\pi}\mid\mathcal{G})=\prod_{t=2}^{n}p_{\bm{\bm{\theta}}}(\pi_{t}\mid\mathcal{G},\pi_{1:t-1}),$ (16)
where the objective function $f(\bm{\pi}|\mathcal{G})$ represents the total reward (e.g., the negative value of tour length) of instance $\mathcal{G}$ given a specific solution $\bm{\pi}$.
Appendix F Adaptation Attention Free Module
Following Zhou et al. (2024a), we implement $\mathrm{Attention}(\cdot)$ using an adaptation attention free module (AAFM) to enhance geometric pattern recognition for routing problems. Given the input $X$, AAFM first transforms it into $Q$, $K$, and $V$ through corresponding linear projection operations:
$Q=XW^{Q},\quad K=XW^{K},\quad V=XW^{V},$ (17)
where $W^{Q}$, $W^{K}$, and $W^{V}$ are learnable matrices. The AAFM computation is then expressed as:
$\mathrm{Attention}(Q,K,V,A)=\sigma(Q)\odot\frac{\exp(A)(\exp(K)\odot V)}{\exp(A)\exp(K)},$ (18)
where $\sigma$ denotes the sigmoid function, $\odot$ represents the element-wise product, and $A=\{a_{ij}\}$ denotes the pair-wise adaptation bias. For distance matrices, the corresponding adaptation bias $f\left(\alpha,N,d_{ij}\right)=-\alpha\cdot\log_{2}{N}\cdot d_{i,j}$. Here, $N$ denotes the total number of nodes (i.e., problem size), and $\alpha$ is generated via a lightweight network $\mathrm{BIAS}(\bm{\lambda})$ in our paper (see Equation 6). Notably, for the relation matrix, we remove the scale $N$ in the calculation of adaptation bias because the relation is scale-independent.
Compared to multi-head attention (MHA) (Vaswani et al., 2017), AAFM enables the model to explicitly capture instance-specific knowledge by updating pair-wise adaptation biases while exhibiting lower computational overhead. Further details are provided in the related work section mentioned above.
Appendix G LLM-driven Constraint Satisfaction Mechanism
To operationalize the UDR framework for a vast spectrum of routing problems, we introduce an automated constraint satisfaction mechanism based on LLMs. This mechanism obviates the need for manual, problem-specific masks by automatically designing executable mask generators. This approach follows the generate-and-validate pipeline, transforming constraint specifications into reliable code.
G.1 Mask Generator Design
This process begins by prompting an LLM with a prompt that comprises two key elements: 1) a natural language task description outlining the constraint logic, and 2) a standardized Python code template that specifies the function signature, inputs, and outputs. Here is an example for generating a mask for the TSP that enforces the fundamental visit-once constraint:
Task Description: Given the current node, visited nodes, and the number of nodes, design a function to mask visited nodes.
Code Template:
1import torch
2
3def mask_visited_nodes(current_node: torch.Tensor, visited_nodes: torch.Tensor, num_nodes: int) -> torch.Tensor:
4 """
5 Design a novel algorithm to mask the visited nodes in each step.
6 The operation processes batches in parallel, so no loops over the batch dimension are allowed-inputs must use batch-first shapes.
7
8 Args:
9 current_node: ID of the current node. Shape: [batch_size, 1]
10 visited_nodes: Array of IDs of visited nodes. Shape: [batch_size, num_visited_nodes]
11 num_nodes: The number of nodes.
12
13 Return:
14 mask: Boolean mask indicating valid nodes (True = selectable, False = masked). Shape: [batch_size, num_nodes]
15 """
16 return mask
The LLM completes the function body with the prompt guidance. The candidate mask generator is an executable Python snippet implementing the described logic:
1import torch
2
3def mask_visited_nodes(current_node: torch.Tensor, visited_nodes: torch.Tensor, num_nodes: int) -> torch.Tensor:
4 """
5 Design a novel algorithm to mask the visited nodes in each step.
6 The operation processes batches in parallel, so no loops over the batch dimension are allowed-inputs must use batch-first shapes.
7
8 Args:
9 current_node: ID of the current node. Shape: [batch_size, 1]
10 visited_nodes: Array of IDs of visited nodes. Shape: [batch_size, num_visited_nodes]
11 num_nodes: The number of nodes.
12
13 Return:
14 mask: Boolean mask indicating valid nodes (True = selectable, False = masked). Shape: [batch_size, num_nodes]
15 """
16 mask = torch.ones(current_node.size(0), num_nodes, dtype=torch.bool)
17 mask.scatter_(1, visited_nodes, 0)
18 mask.scatter_(1, current_node, 0)
19 return mask
G.2 Automated Validation
Each candidate generator is then validated. We execute it within the URS solver or a simple simulated solver on a set of problem instances and use a problem-specific checker function to verify if the resulting solutions are constraint-compliant. For the TSP example, the checker ensures that every node is visited exactly once:
1import torch
2
3def checker(solutions: torch.Tensor, problem_size: int, num_instances: int) -> float:
4 """
5 Check the validity of TSP solutions that use the generated masks.
6
7 Args:
8 solutions: A batch of generated solutions to be validated. Each row represents a single solution sequence. Shape: [batch_size, problem_size].
9 problem_size: The problem size.
10
11 Returns:
12 valid_rate: The validity rate, representing the fraction of valid solutions in the batch.
13 """
14 num_instances = solutions.shape[0]
15 sorted, _ = torch.sort(solutions, dim=1)
16 valid = (sorted == torch.arange(problem_size)[None, :].expand(num_instances, problem_size)).all(1)
17 valid_rate = torch.mean(valid.float()).item()
18 return valid_rate
G.3 AHD Framework Integration
The mask generator design process can be formulated as an optimization problem aimed at maximizing the solution validity rate across test instances. This allows us to leverage sophisticated LLM-based automated algorithm design frameworks, such as EoH (Liu et al., 2024b), to improve the efficiency, especially for the complex VRP variants. The search terminates as soon as a fully reliable generator is found, making the process significantly more efficient than general optimization tasks. This LLM-driven synthesis, combined with UDR, empowers URS with true zero-shot generalization capabilities while rigorously upholding solution feasibility across an unprecedented range of routing problems.
Appendix H Training and Model Settings
The detailed information about the hyperparameter settings can be found in Table 9. The training of URS is conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB memory). The training procedure requires only about $95$ hours ($\approx 4$ days) in total (500 epochs at approximately $12$ minutes each) to obtain a unified parameter set capable of addressing a broad class of routing problems.
Table 9: The hyperparameter settings of URS.
Hyperparameter Value
Optimizer AdamW
Clipping parameter $\zeta$ 50
Initial learning rate $10^{-4}$
Epochs of learning rate decay [451]
Factor of learning rate decay 0.1
Weight decay $10^{-6}$
The number of encoder layers $L$ $12$
Embedding dimension $d$ $128$
Feed forward dimension $512$
Dimension $d_{h}$ in Equation 12 $256$
Training scale $100$
Batches of each epoch $2,000$
Batch size $128$
Training Epochs $500$
Appendix I Detailed Experimental Results
To comprehensively evaluate zero-shot generalization of URS, we test its performance on a diverse set of 96 unseen problems. This set is designed to cover both established and novel challenges, including 44 well-studied VRP variants from the NCO literature and 52 newly introduced variants that have not been explored before. The detailed results are presented in Table 10 and Table 11. On widely-studied constraints (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions. Notably, URS merely selects the next node from a given feasible candidates without any additional domain knowledge, making its performance especially meaningful. Benefiting from our UDR, URS can directly address previously unexplored variants, which compound asymmetry and PD relations. Without any fine-tuning, URS still constructs high-quality solutions for them, strengthening the broad problem generality and deployment efficiency.
Table 10: Zero-shot generalization performance across $44$ VRP variants that have been explored in current literature (Liu et al., 2024a; Berto et al., 2024) ($N=100$).
Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 13.95% 14s 17.36% 14s 13.38% 14s 2.29% 15s 10.46% 16s
MTPOMO 13.71% 7s 16.94% 8s 13.44% 8s 3.26% 9s 11.35% 10s
ReLD-MTL 13.57% 9s 14.83% 9s 12.96% 10s 2.27% 11s 10.80% 12s
[…]
Method CVRPBPTW OCVRPBPL OCVRPBPLTW CVRPL CVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 16m 0.00% 3.5h
MVMoE 10.85% 15s 16.88% 14s 1.84% 16s 0.26% 12s 1.35% 12s
MTPOMO 11.53% 9s 16.39% 8s 2.77% 10s 0.48% 6s 1.79% 6s
[…]
Method OCVRPB CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h
MVMoE 7.09% 12s 1.47% 16s 9.95% 15s 7.33% 16s 3.15% 13s
MTPOMO 7.34% 5s 1.92% 9s 10.45% 6s 7.75% 8s 3.44% 6s
[…]
URS 9.35% 6s 2.67% 9s 13.77% 7s 9.18% 8s 3.22% 7s
Method CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW MDCVRPBP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 6.6m
MVMoE 7.08% 15s 7.12% 12s 3.90% 15s 10.01% 15s 64.91% 33s
MTPOMO 7.41% 7s 7.34% 7s 4.37% 8s 10.50% 7s 45.66% 24s
[…]
Method MDOCVRPBP MDCVRPBPL MDOCVRPBPTW MDCVRPBPLTW MDCVRPBPTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 52.16% 33s 64.79% 36s 46.18% 38s 54.23% 42s 54.76% 39s
MTPOMO 40.15% 24s 45.84% 26s 34.19% 29s 40.26% 32s 40.32% 30s
ReLD-MTL 34.53% 27s 39.06% 29s 34.08% 34s 36.59% 37s 37.16% 35s
[…]
Method MDOCVRPBPL MDOCVRPBPLTW MDCVRP MDCVRPTW MDOCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 51.72% 34s 45.23% 40s 31.90% 31s 53.64% 41s 29.08% 31s
MTPOMO 39.34% 27s 33.37% 32s 23.75% 23s 36.10% 31s 27.46% 23s
[…]
Method MDCVRPL MDCVRPB MDOCVRPTW MDOCVRPB MDCVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 33.42% 34s 21.70% 26s 51.57% 40s 28.21% 26s 20.35% 29s
MTPOMO 24.32% 25s 12.89% 19s 35.71% 29s 25.06% 20s 11.10% 20s
[…]
Method MDCVRPLTW MDOCVRPBTW MDCVRPBLTW MDOCVRPL MDCVRPBTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 55.55% 44s 62.47% 35s 63.18% 37s 29.33% 33s 65.90% 35s
MTPOMO 37.34% 32s 45.55% 25s 45.07% 27s 27.21% 25s 45.94% 25s
[…]
Method MDOCVRPBL MDOCVRPLTW MDOCVRPBLTW SPCTSP -
Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 1.6h -
MVMoE 28.70% 28s 51.09% 43s 62.10% 37s - -
MTPOMO 25.01% 21s 35.39% 31s 45.48% 27s - -
ReLD-MTL 18.50% 22s 33.26% 35s 43.27% 30s - -
AM - - - 2.69% 5s -
[…]
Table 11: Zero-shot generalization performance across $52$ VRP variants that have not yet been explored in the current literature ($N=100$).
Method ACVRPL ACVRPBL ACVRPLTW ACVRPBLTW AOCVRPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPBL AOCVRPLTW AOCVRPBLTW ACVRPBPL ACVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPBPL AOCVRPBPLTW ACVRPB ACVRPTW AOCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPTW AOCVRPB AOCVRPBTW ACVRPBTW ACVRPBP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AOCVRPBP AOCVRPBPTW ACVRPBPTW APDTSP AMDCVRPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDCVRPBL AMDCVRPLTW AMDCVRPBLTW AMDOCVRPL AMDOCVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPLTW AMDOCVRPBLTW AMDCVRPBPL AMDCVRPBPLTW AMDOCVRPBPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPBPLTW AMDCVRP AMDCVRPTW AMDOCVRP AMDCVRPB
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPTW AMDOCVRPB AMDOCVRPBTW AMDCVRPBTW AMDCVRPBP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method AMDOCVRPBP AMDOCVRPBPTW AMDCVRPBPTW PDCVRP OPDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
[…]
Method APDCVRP AOPDCVRP - - -
Gap Time Gap Time
Oracle Solver 0.00% 6.6m 0.00% 6.6m - - -
[…]
Appendix J Results on Benchmark Dataset
Our test results on Benchmark datasets are presented in Table 12. All models are trained on instances of size $N=100$. Some results for comparative models are sourced from MVMoE (Zhou et al., 2024b) and RouteFinder (Berto et al., 2024). URS maintains its position as the best overall performance across instances of varying scales. Notably, RL-based URS surpasses the SL-based representative specialized model LEHD (8.678% vs. 12.836%), underscoring its practical applicability in real-world scenarios.
Table 12: Results on large-scale CVRPLIB instances (Set-X) (Uchoa et al., 2017) ($500\leq N\leq 1000$). All models are trained on instances of size $N=100$.
Set-X POMO LEHD MTPOMO MVMoE/4E MVMoE/4E-L RF-MVMoE RF-TE URS
Instance Opt. Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap
[…]
Appendix K Ablation Study
In this section, we conduct a detailed ablation study and analysis to demonstrate the effectiveness and robustness of URS. Please note that, unless stated otherwise, the results presented in the ablation study reflect the best result from multiple trajectories, and we employ instance augmentation to improve performance in the ablation study. We adopt the widely used 100-node setting as our primary evaluation scenario for URS.
K.1 Effects of Node-type Indicator
Table 13: The ablation study of the node-type indicator on trained VRPs
Method ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
[…]
To evaluate the effectiveness of node-type indicator $\bm{\xi}$, we conduct an ablation study about including and excluding $\bm{\xi}$. Table 13 shows that adding the node-type indicator reduces the average optimality gap from 3.78% to 2.74%, with large gains on the routing problem with relatively rich instance node roles, such as PDTSP, where the coexistence of a depot and paired pickup–delivery nodes (with inherent precedence coupling) allows the explicit node-role encoding to further boost performance. Meanwhile, URS maintains a competitive performance in addressing the problem of single-node roles.
K.2 Effects of Problem State
Table 14: Performance impact of removing the problem state feature $C_{t}$
Method ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
[…]
As described in Appendix E.3, the majority of VRP variants are already solved without supplying an additional problem state $C_{t}$. To test whether the remaining use of $C_{t}$ matters, we conduct an ablation in which we eliminate it entirely, enforcing $C_{t}\equiv 0$ for all problems. The results in Table 14 show that URS maintains its competitive performance even in the absence of $C_{t}$. This indicates that explicit constraint-state features are not essential for the effectiveness of URS. Interestingly, on (O)CVRPTW instances, the model without any explicit constraint features outperforms the version that supplies only capacity information. We guess that capacity and time windows information are equally important, so enforcing only capacity creates a bias toward load considerations.
K.3 Effects of Mixed Bias Module
Table 15: Ablation of prior components in the Mixed Bias Module (MBM).
$\bm{D}$ $\bm{D^{\mathrm{T}}}$ $\bm{R}$ ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
[…]
To further assess the effectiveness of MBM, we conduct an ablation study that incorporates different combinations of problem-specific priors into the MBM. Owing to its flexibility, the MBM permits imposing multiple priors on the same shared attention layer. The results in Appendix K.3 underscore the complementary roles of the three prior components: (1) The transposed distance matrix $\bm{D^{\mathrm{T}}}$ captures directional asymmetry absent from the raw distance matrix, and its inclusion markedly stabilizes optimization on asymmetric variants; (2) Removing the pickup–delivery relation matrix $\bm{R}$ degrades performance on PDTSPs, confirming that explicit relational coupling is indispensable for these instances; (3) When all priors are fused, MBM can simultaneously learn the geometric and relational biases inherent in various problems, yielding the best overall performance.
K.4 Effects of Different Decoder Architectures
Table 16: Effect of different decoder architectures on URS cross‑problem performance
Decoder ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
URS-POMO 4.83% 0.74% 0.72% 1.29% 5.45% 5.09% 1.99% 6.42% 1.83% 3.62% 5.31% 3.39%
URS-ReLD 6.48% 0.90% 0.86% 1.51% 5.81% 6.51% 2.33% 6.56% 2.33% 3.62% 5.64% 3.87%
[…]
We adopt an AM (Kool et al., 2019) as our basic encoder-decoder model. Multiple decoder refinements have been proposed around AM (Kwon et al., 2020; Huang et al., 2025; Zhou et al., 2024a). To assess how decoder structure impacts multi-task performance, we additionally train two URS variants whose decoders follow POMO and ReLD, denoted URS-POMO and URS-ReLD. For stronger and fairer baselines, we apply Distance-aware Attention Reshaping (DAR) (Wang et al., 2025b) to both, and all decoder parameters are generated by a hypernetwork conditioned on the problem representation $\bm{\lambda}$ (see Appendix E.2). As shown in Table 16, URS-POMO and URS-ReLD still exhibit strong cross-problem robustness, while replacing the decoder with ICAM (Zhou et al., 2024a) yields a further overall improvement.
Appendix L Licenses for Used Resources
Table 17: List of licenses for the codes and datasets we used in this work.
Resource Type Link License
HGS-PyVRP (Wouda et al., 2024) Code https://github.com/PyVRP/PyVRP MIT License
LKH3 (Helsgaun, 2017) Code http://webhotel4.ruc.dk/~keld/research/LKH-3/ Available for academic research use
OR-Tools (Perron & Furnon, 2023) Code https://github.com/google/or-tools Apache-2.0 License
POMO (Kwon et al., 2020) Code https://github.com/yd-kwon/POMO/tree/master/NEW_py_ver MIT License
Sym-NCO (Kim et al., 2022) Code https://github.com/alstn12088/Sym-NCO Available for any non-commercial use
LEHD (Luo et al., 2023) Code https://github.com/CIAM-Group/NCO_code/tree/main/single_objective/LEHD Available for any non-commercial use
BQ (Drakulic et al., 2023) Code https://github.com/naver/bq-nco CC BY-NC-SA 4.0 license
ICAM (Zhou et al., 2024a) Code https://github.com/CIAM-Group/ICAM MIT License
MatNet (Kwon et al., 2021) Code https://github.com/yd-kwon/MatNet/tree/main MIT License
Heter-AM (Li et al., 2021) Code https://github.com/jingwenli0312/Heterogeneous-Attentions-PDP-DRL MIT License
MTPOMO (Liu et al., 2024a) Code https://github.com/FeiLiu36/MTNCO MIT License
MVMoE (Zhou et al., 2024b) Code https://github.com/RoyalSkye/Routing-MVMoE MIT License
RouteFinder (Berto et al., 2024) Code https://github.com/ai4co/routefinder MIT License
CaDA (Li et al., 2025a) Code https://github.com/CIAM-Group/CaDA MIT License
ReLD (Huang et al., 2025) Code https://github.com/ziweileonhuang/reld-nco MIT License
GOAL (Drakulic et al., 2025) Code https://github.com/naver/goal-co Available for any non-commercial use
CVRPLIB Set-X (Uchoa et al., 2017) Dataset http://vrp.galgos.inf.puc-rio.br/index.php/en/ Available for academic research use
We list the used existing codes and datasets in Table 17, and all of them are open-sourced resources for academic usage.
Appendix M LLM usage
During the preparation of this manuscript, we utilized a Large Language Model (LLM) as a writing assistance tool. The primary purpose of its use was to improve the linguistic quality of the paper. Specifically, the LLM is used to polish the writing for enhanced clarity and fluency and to identify and correct grammatical, spelling, and punctuation errors.
We confirm that all core research ideas, study design, data analysis, and conclusions are independently developed by the authors. The LLM does not contribute to the research ideation, methodology, or substantive academic content of the paper. Its function was strictly confined to language editing and proofreading to ensure the manuscript’s professionalism and readability.