### v2

# URS: A Unified Neural Routing Solver for Cross-Problem Zero-Shot Generalization

Abstract
Multi-task neural routing solvers have emerged as a promising paradigm for their ability to solve multiple vehicle routing problems (VRPs) using a single model. However, existing neural solvers typically rely on predefined problem constraints or require per-problem fine-tuning, which substantially limits their zero-shot generalization ability to unseen VRP variants. To address this critical bottleneck, we propose URS, a unified neural routing solver that achieves zero-shot generalization across a wide range of unseen VRPs with a single model. We propose a unified data representation (UDR) that replaces problem enumeration with data unification, thereby broadening the problem coverage and reducing reliance on domain expertise. In addition, we introduce a Mixed Bias Module (MBM) during encoding to improve node embeddings, which efficiently captures multiple priors inherent to various problems. On top of the UDR, we develop a problem-conditioned parameter generator to further improve zero-shot generalization. Extensive experiments show that URS consistently produces high-quality solutions for 110 VRP variants (including 99 unseen variants) while demonstrating impressive scalability to large-scale instances with up to 7000 nodes. To the best of our knowledge, URS is the first neural solver to handle over 100 VRP variants with a single model. Our code is available at https://github.com/CIAM-Group/URS.
Machine Learning, ICML
1 Introduction
The Vehicle Routing Problem (VRP) is an essential class of combinatorial optimization problems (COPs) with extensive applications in logistics and supply chain management (Tiwari and Sharma, 2023; Sar and Ghadimi, 2023). Solving VRPs efficiently is challenging due to their NP-hard nature. While exact solvers can produce optimal solutions, they often become computationally prohibitive for real-world scenarios. In recent decades, classical heuristic solvers (Helsgaun, 2017; Vidal, 2022) have achieved impressive performance within acceptable timeframes. Nevertheless, these methods require considerable domain expertise to design specialist rules for each routing problem. Given the growing diversity of VRP variants in real-world applications, manually crafting tailored rules for every case has become impractical.
In recent years, neural combinatorial optimization (NCO) methods have attracted substantial attention for their potential to reduce reliance on handcrafted rules while maintaining competitive solution quality (Bengio et al., 2021; Ba et al., 2026). This has led to the development of many high-performing neural routing solvers that automatically learn implicit problem-specific rules from data under different training paradigms such as supervised learning (SL) (Drakulic et al., 2023; Luo et al., 2023, 2025b), reinforcement learning (RL) (Kool et al., 2019; Zhou et al., 2026), and self-improved learning (SIL) (Luo et al., 2025a; Pirnay and Grimm, 2024). Although these solvers have demonstrated impressive performance on specific problems (Huang et al., 2025; Zhou et al., 2025), they typically require architectural customization and per-problem retraining to accommodate the distinct constraints and features of different VRP variants, thereby increasing overall training costs and hindering practical deployment to new problems.
Table 1: Comparison between our URS and existing neural solvers with multi-task learning. Note that ”#\#VRP Variants” and ”Generalizable Scale” refer to the total number of tested problems and the maximum scale reported in the original papers, respectively.
Multi-task Neural Routing Solver Training #\#VRP Variants Symmetric Asymmetric Pickup and Delivery Zero-shot Training Generalizable Remarks
Paradigm VRPs VRPs Problems Generalization Scale Scale
MTPOMO (Liu et al., 2024) RL 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
MVMoE (Zhou et al., 2024) RL 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
RouteFinder (Berto et al., 2025) RL 48 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
CaDA (Li et al., 2025a) RL 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $200$ Constraint Combination
MTL-KD (Zheng et al., 2025) RL+KD 16 $\checkmark$ $\times$ $\times$ $\checkmark$ $100$ $1000$ Constraint Combination
TSP-FT (Lin et al., 2024) RL 4 $\checkmark$ $\times$ $\times$ $\times$ $100$ $1000$ Adapter-based Fine-tuning
MTL-MAB (Wang et al., 2025a) RL 3 $\checkmark$ $\times$ $\times$ $\times$ $100$ $1000$ Adapter-based Fine-tuning
GOAL (Drakulic et al., 2025) SL 10 $\checkmark$ $\checkmark$ $\times$ $\times$ $100$ $1000$ Adapter-based Fine-tuning
URS (Ours) RL 110 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 100 7000 Unified Data Representation
To address the challenge of cross-problem generalization, growing attention has been directed toward multi-task learning capable of handling diverse routing problems. As summarized in Table 1, existing efforts largely fall into two categories: 1) constraint combination-based multi-task learning and 2) adapter-based fine-tuning. In the first approach, VRP variants are treated as different combinations of constraint attributes, and a unified model is trained across these combinations to enable knowledge sharing for problems with seen constraints (Liu et al., 2024; Zhou et al., 2024; Li et al., 2025a; Berto et al., 2025; Zheng et al., 2025). The second approach builds a shared model backbone for all problems and incorporates problem-specific input/output adapters, thereby reducing retraining costs (Lin et al., 2024; Drakulic et al., 2025; Wang et al., 2025a).
Despite the above advancements, existing methods still fall short in cross-problem zero-shot generalization. The problem coverage of constraint combination-based approaches is inherently bounded by their manually specified constraint sets, whereas adapter-based approaches require additional fine-tuning and thus cannot perform zero-shot generalization. Furthermore, both strategies fundamentally rely on the explicit problem enumeration through a predefined set of problem tags. This practice is challenging because the constraint space of VRP variants is open-ended and compositional. More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which NCO always aims to avoid. A detailed review of related work on single-task and multi-task learning is provided in Appendix A.
In this paper, we propose a powerful Unified Neural Routing Solver (URS) to improve the cross-problem zero-shot generalization ability for NCO methods. Our contributions can be summarized as follows: (1) We propose a unified data representation (UDR) that replaces existing problem enumeration with data unification, significantly broadening problem coverage while mitigating reliance on domain-specific expertise; (2) We introduce a Mixed Bias Module (MBM) during encoding to improve node embeddings, which efficiently captures multiple priors inherent to various problems while reducing architectural redundancy; (3) On top of the UDR, we develop a problem-conditioned parameter generator to further improve zero-shot generalization; (4) Extensive experiments on 110 VRP variants show that URS not only achieves competitive performance against specialist neural solvers on seen variants but also exhibits strong zero-shot generalization across 99 unseen variants. Notably, URS demonstrates impressive scalability to large-scale instances with up to 7000 nodes. To the best of our knowledge, URS is the first neural solver to efficiently solve over 100 VRP variants with a single model, without retraining or fine-tuning.
2 Preliminaries
In this section, we first introduce the definition of VRP, then provide an overview of recent constructive neural solvers for solution generation (Kool et al., 2019; Kwon et al., 2020).
Vehicle Routing Problems
Consider a VRP instance with an optional depot indexed by 0 and $n$ customers indexed by $\{1,2,\ldots,n\}$, such as in the Capacitated VRP (CVRP). The instance can be represented as a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where the node set $\mathcal{V}=\{v_{i}\}_{i=0}^{n}$ has a total size of $|\mathcal{V}|=1+n$ unless the variant has no depot (e.g., the Traveling Salesman Problem), and the edge set is $\mathcal{E}=\{e(v_{i},v_{j})\ |\ v_{i},v_{j}\in\mathcal{V},v_{i}\neq v_{j}\}$. The travel cost between nodes is given by a distance matrix $\bm{D}=\{d_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$. In VRP, each node $v_{i}\in\mathcal{V}$ includes node coordinates $\{x_{i},y_{i}\}$ when available and problem-specific attributes (e.g., demands in CVRP). Formally, a feasible solution is denoted by a node permutation $\bm{\pi}=(\pi_{1},\pi_{2},\ldots,\pi_{m})$ (Toth and Vigo, 2002) (i.e., a finite sequence) that represents a set of valid routes. Each route is traversed by a single vehicle originating and terminating at its own depot or starting node, while strictly satisfying all constraints. Let $\Omega$ denote the feasible sequence set, given an objective function $f(\bm{\pi}|\mathcal{G})$, we aim to search for a sequence $\bm{\pi}^{*}$ with the maximum objective, i.e., $\bm{\pi}^{*}=\arg\max_{\bm{\pi}\in\Omega}f(\pi|\mathcal{G})$. In most VRPs, $f(\bm{\pi}|\mathcal{G})$ can be defined as the negative value of the total distance of $\bm{\pi}$.
In this paper, we consider 110 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); and (10) Pickup and Delivery (PD). Detailed definitions of these constraints are provided in Appendix B.
Figure 1: The pipeline of our URS for solving 110 VRP variants using a single model without any fine-tuning, illustrated on APDTSP.
Constructive Neural Routing Solver
Most constructive NCO methods employ an encoder-decoder architecture for solution construction (Luo et al., 2023; Kwon et al., 2020). Let learnable model parameters be denoted as $\bm{\theta}=\{\bm{\theta}_{enc},\bm{\theta}_{dec}\}$. Without loss of generality, we present the prevailing autoregressive construction pipeline using AM (Kool et al., 2019). Given an instance $\mathcal{G}$, raw node features are first mapped by a linear projection to initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$. They are then processed by an encoder $\bm{\theta}_{enc}$ with $L$ attention layers to produce refined node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. At each decoding step $t$, the decoder $\bm{\theta}_{dec}$ sequentially selects a node to append to the current partial solution $\pi_{1:t-1}=(\pi_{1},\pi_{2},\ldots,\pi_{t-1})$, where $\pi_{1},\pi_{t-1}\in\mathcal{V}$ are the first and last visited node, respectively. This sequential process continues until a complete solution is formed. To guarantee the feasibility of generated solutions, a masking function $\mathcal{M}_{t}$ sets the selection probabilities of the following nodes to -$\infty$ during the construction process: nodes that (1) are visited or (2) violate problem-specific constraints (e.g., capacity).
3 Methodology
In this section, as illustrated in Figure 1, we propose a powerful Unified Neural Routing Solver (URS), which significantly improves cross-problem zero-shot generalization for NCO methods. Its key components are elaborated below.
3.1 Unified Data Representation
Existing multi-task neural solvers (Zhou et al., 2024; Drakulic et al., 2025) fundamentally rely on explicit problem enumeration via a predefined set of problem tags, which is challenging because the constraint space of VRPs is open-ended and compositional. More critically, creating and maintaining such a problem taxonomy requires considerable domain expertise, which NCO always aims to avoid.
From a data perspective, despite the substantial diversity in constraints, their instances share a common structural representation. For existing constraint combination-based methods (Liu et al., 2024; Zhou et al., 2024), their evaluated problems can be formulated as CVRPTW instances. Leveraging this structural commonality, we propose a Unified Data Representation (UDR) $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$. By replacing explicit problem enumeration with data unification, our UDR broadens problem coverage and reduces reliance on domain expertise. Notably, the inputs to existing multi-task neural solvers based on either problem-specific adapters (Drakulic et al., 2025) or constraint combinations (Zhou et al., 2024) can be viewed as special subsets of our proposed UDR.
Instead of relying on a predefined problem taxonomy with a discrete set of problem tags, UDR decouples instance representation from constraint definitions and operates in a unified space characterized by both continuous and discrete features. It allows a single neural solver to address a much larger, open-ended space of VRP variants, as new problems can be seamlessly incorporated into this unified representation. Problem-specific constraints are delegated to the model-agnostic masking function $\mathcal{M}_{t}$. For example, for all CVRP variants, only the remaining load is explicitly provided as input, while diverse additional constraints (e.g., time windows) are enforced implicitly by masking infeasible candidate nodes during solution construction, rather than being fully enumerated in the decoder’s input. Thus, UDR enables a single model to handle various VRP variants while significantly reducing reliance on domain expertise. Specifically, for an instance $\mathcal{G}$, each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a positional identifier, unified attribute set, and node-type indicator, respectively.
Positional Identifier ($\bm{\rho}_{i}$)
We define a unified positional identifier for each node as $\bm{\rho}_{i}=\{\eta_{i},x_{i},y_{i}\}\in[0,1]^{3}$. The $\eta_{i}\sim\mathrm{Uniform}(0,1)\in\mathbb{R}^{1}$ is a sampled scalar, which is used to address asymmetric problems, following  Drakulic et al. (2023). For symmetric cases, we simply set $\eta_{i}=0$.
Unified Attribute Set ($\bm{\omega}_{i}$)
We define the unified attribute set as $\bm{\omega}_{i}=\{\delta_{i},\epsilon_{i},\mu_{i},\ e_{i},\ l_{i},\ s_{i}\}$, where they correspond to demand, prize, penalty, earliest arrival time, latest arrival time, and service time, respectively. The unification empowers the model to learn the relative importance of individual attributes in a shared embedding space. It also supports effortless attribute extension or ablation via simple zero-filling, eliminating the need to change the architecture.
Node-type Indicator ($\bm{\xi}_{i}$)
We provide a 5-dimensional binary vector $\bm{\xi}_{i}\in\{0,1\}^{5}$ for each node, encoding: depot, pickup node, delivery node, node in sub-routes, node in open route. The nodes in sub-routes and open routes mean the solution $\pi$ can have sub-routes or be an open route. We introduce explicit structural roles to enable the model to generalize over diverse variants. Further ablation analysis concerning $\bm{\xi}_{i}$ is provided in Appendix K.1.
Multi-hot Problem Representation ($\bm{\lambda}_{i}$)
Through the UDR, each problem instance activates only a specific subset of the unified feature space. For any instance $\mathcal{G}_{i}$ drawn from different problems, a corresponding multi-hot problem representation $\bm{\lambda}_{i}$ is obtained by indicating active (non-zero) feature slots with a value of $1$. This representation captures only the presence of features, in contrast to their raw values, and it subsequently provides conditional guidance to the position bias weight (see Equation 6) and the adaptive decoder parameters $\bm{\theta}_{dec}(\bm{\lambda})$ (see Equation 9 and Equation 10), thereby helping the model in distinguishing problem variants and refining its node selection process.
A more detailed description of $\bm{U}$ and $\bm{\lambda}$ is provided in Appendix C and Appendix D, respectively.
3.2 Model Architecture
As shown in Figure 1, we adopt AM (Kool et al., 2019) as our basic model. While UDR provides cross-problem consistency, different problems have diverse geometric priors. To efficiently learn the multiple priors inherent in various problems and obtain better cross-problem zero-shot generalization, we propose two key enhancements: (1) a Mixed Bias Module (MBM) that captures multiple priors inherent to various problems, and (2) problem-conditioned parameter generators conditioned on the UDR. Their implementations are detailed below.
Embedding Layer
Given an instance $\mathcal{G}$, for each $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, we project it into $d$-dimensional embeddings through three separate linear transformations:
$\mathbf{h}_{i}^{(0)}=\bm{\rho}_{i}W_{\bm{\rho}}+\bm{\omega}_{i}W_{\bm{\omega}}+\bm{\xi}_{i}W_{\bm{\xi}},$ (1)
where $W_{\bm{\rho}}\in\mathbb{R}^{3\times d},\ W_{\bm{\omega}}\in\mathbb{R}^{6\times d},\ W_{\bm{\xi}}\in\mathbb{R}^{5\times d}$ are learnable matrices. Then we obtain a set of initial embeddings $H^{(0)}=\{\mathbf{h}_{i}^{(0)}\}_{i=0}^{n}$ for all nodes in instance $\mathcal{G}$. This embedding is the initial input of encoder $\bm{\theta}_{enc}$. After passing through $L$ stacked attention layers, $H^{(0)}$ is transformed into advanced node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. The detailed encoding process is provided in Appendix E.1.
Figure 2: The structure of the proposed Mixed Bias Module, explained on a 44-node APDTSP instance. We replace plain attention with MBM during encoding to improve node embeddings, which effectively captures multiple priors inherent to various problems while reducing architectural redundancy.
Mixed Bias Module
Given the distinct priors across VRP variants, efficiently capturing the intrinsic geometric bias is pivotal for enhancing the cross-problem generalization of RL-based multi-task neural solvers. Current leading RL-based models generally adopt a heavy encoder and light decoder architecture, where the quality of the encoder-generated node embeddings plays a key role in overall performance. While prior studies like MatNet (Kwon et al., 2021) have attempted to handle asymmetry by integrating directional distance biases via parallel layers, such designs incur high model complexity and lack universality. Consequently, effectively encoding multiple inherent biases, such as symmetric and asymmetric distances, and relational dependencies, remains challenging, hindering the acquisition of high-quality solutions across a wide range of problems.
To address the limitation, we propose the mixed bias module (MBM) to replace the vanilla attention mechanism (Vaswani et al., 2017), which enhances cross-problem generalization while reducing architectural redundancy. As shown in Figure 2, MBM can efficiently capture three problem-specific prior matrices via three separate attention calculations: (1) an outgoing distance matrix $\bm{D}$; (2) an incoming distance matrix $\bm{D}^{\mathrm{T}}$; and (3) an optional relation matrix $\bm{R}=\{r_{ij}|\ \forall\ i,j\in 0,\ldots,n\}$, where $r_{ij}=0$ if a predefined relation exists (e.g., pickup-delivery pairing) and 1 otherwise. This ensures that smaller values are preferred, aligning with the distance metric, where smaller values correspond to a larger selection bias. For the $\ell$-th layer, the MBM output $\hat{\mathbf{h}}_{i}^{(\ell)}$ can be expressed as:
$\bar{\mathbf{h}}_{i}^{(0)}=\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,N,\bm{D}_{i}\right)),$ (2)
$\bar{\mathbf{h}}_{i}^{(1)}=\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,N,\bm{D}_{i}^{\mathrm{T}}\right)),\\$ (3)
$\bar{\mathbf{h}}_{i}^{(2)}=\begin{cases}\mathrm{Attention}(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},f\left(\alpha,\bm{R}_{i}\right))&\text{if}\ \bm{R}\neq\emptyset\\ \bm{0}&\text{otherwise},\end{cases}$ (4)
$\hat{\mathbf{h}}_{i}^{(\ell)}=\left[\bar{\mathbf{h}}_{i}^{(0)},\bar{\mathbf{h}}_{i}^{(1)},\bar{\mathbf{h}}_{i}^{(2)}\right]W^{O},$ (5)
where $\cdot$$\cdot$] denotes the horizontal concatenation operator, $W^{O}\in\mathbb{R}^{(3\times d)\times d}$ is a learnable matrix. For $\mathrm{Attention}(\cdot)$, we adopt the attention mechanism introduced by Zhou et al. (2026), which boosts the perception of diverse geometric patterns through an adaptation function $f(\alpha,N,d_{ij})$ with bias weight $\alpha$ (see Appendix F for implementation details). Note that since the relation $r_{ij}$ is scale-independent, we remove the scale $N$ in the adaptation function for $\bm{R}$. If $\bm{R}$ is absent, we substitute a zero vector for $\bar{\mathbf{h}}_{i}^{(2)}$. For a comparison between our MBM and existing attentions (Vaswani et al., 2017; Kwon et al., 2021), please see Appendix G.
Problem-Conditioned Parameter Generation
Using a single static set of parameters to solve 110 VRP variants is a significant challenge. To mitigate this, we introduce a problem-conditioned mechanism based on problem representation $\bm{\lambda}$, which consists of two components: (1) $\mathrm{BIAS}(\bm{\lambda})$ adjusts the degree to which different priors are integrated into the attention; and (2) $\mathrm{WEIGHT}(\bm{\lambda})$ generates decoder parameters for each problem.
As shown in Figure 3, unlike existing adapter-based fine-tuning methods (Lin et al., 2024; Drakulic et al., 2025), we provide a unified decoding entrance and enable URS to adaptively generate decoder parameters for each problem. This conditional weight generation removes problem-specific adapters, enabling zero-shot generalization to unseen problems while maintaining solution quality.
For MBM in Equation 5, we generate $\alpha$ via a lightweight bias network $\mathrm{BIAS}(\bm{\lambda})$:
$\mathrm{BIAS}(\bm{\lambda})=\max\left(1,\ \left(\bm{\lambda}W_{1}+\mathbf{b}_{1}\right)W_{2}+\mathbf{b}_{2}\right),$ (6)
where $W_{1}\in\mathbb{R}^{|\bm{\lambda}|\times d},\ W_{2}\in\mathbb{R}^{d\times 1},\ \mathbf{b}_{1}\in\mathbb{R}^{d},\ \mathbf{b}_{2}\in\mathbb{R}^{1}$ are learnable parameters. We set the minimum bias to $1$, resulting in faster model convergence.
For decoder $\bm{\theta}_{dec}(\bm{\lambda})$, all parameters are generated based on $\bm{\lambda}$. Following Kwon et al. (2020), given the first and last visited node embeddings $\mathbf{h}_{\pi_{1}}^{(L)}$ and $\mathbf{h}_{\pi_{t-1}}^{(L)}$, and an optional constraint state $C_{t}\in\mathbb{R}^{1}$ (e.g., remaining load), the context embedding $\mathbf{h}_{(C)}^{t}$ can be expressed as:
$\mathbf{h}_{(C)}^{t}=\left[\mathbf{h}_{\pi_{1}}^{(L)},\ \mathbf{h}_{\pi_{t-1}}^{(L)},\ C_{t}\right]W_{Q}(\bm{\lambda}),$ (7)
where $\cdot$$\cdot$] is the horizontal concatenation operator, $W_{Q}(\bm{\lambda})\in\mathbb{R}^{(2d+1)\times d}$ is a linear projection matrix conditioned on $\bm{\lambda}$. Missing $C_{t}$ is zero-padded (see Appendix E.3 for details). The new context embedding $\hat{\mathbf{h}}_{(C)}^{t}$ is obtained via $\mathrm{Attention}(\cdot)$ (Zhou et al., 2026) on $\mathbf{h}_{(C)}^{t}$ and $H^{(L)}$:
$K=H^{(L)}W_{K}(\bm{\lambda}),\qquad V=H^{(L)}W_{V}(\bm{\lambda}),$ (8)
$\hat{\mathbf{h}}_{(C)}^{t}=\mathrm{Attention}(\mathbf{h}_{(C)}^{t},K,V,\mathcal{M}_{t},f(\alpha,N,d_{t,i})),$ (9)
where $W_{K}(\bm{\lambda}),W_{V}(\bm{\lambda})\in\mathbb{R}^{d\times d}$ are linear projection matrices conditioned on $\bm{\lambda}$. Finally, we compute probabilities for feasible nodes $p_{\bm{\theta}}(\pi_{t}=i\mid\mathcal{G},\pi_{1:t-1})=\mathrm{softmax}(\mathbf{u})$ based on the improved compatibility (Zhou et al., 2026):
$u^{t}_{i}=\begin{cases}\zeta\cdot\text{tanh}(\frac{\hat{\mathbf{h}}_{(C)}^{t}(\mathbf{h}_{i}^{(L)})^{\mathrm{T}}}{\sqrt{d}}+f(\alpha,N,d_{t,i}))&\text{if}\ i\not\in\{\pi_{1:t-1}\}\\ -\infty&\text{otherwise,}\end{cases}$ (10)
Figure 3: Comparison between our URS and existing adapter-based fine-tuning frameworks.
where $\zeta$ is the clipping parameter. Inspired by Lin et al. (2022), we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ (see Appendix E.2 for implementation details). For $\alpha$ in Equation 9 and Equation 10, we generate them via $\mathrm{BIAS}(\bm{\lambda})$ in Equation 6.
Table 2: Performance comparison on $8$ seen less-explored variants.
Method OP100 PCTSP100 PDTSP100 ACVRP100 CVRPTW100 CVRPB100 OCVRP100 OCVRPTW100
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 1.5m 0.00% 1.2h 0.00% 9.8m 0.00% 6.6m 0.00% 19.6m 0.00% 20.8m 0.00% 5.3m 0.00% 20.8m
Sym-NCO 0.68% 15s 0.14% 16s - - - - - -
BQ 0.63% 7s - - - - - - -
Heter-AM - - 6.61% 5m - - - - -
MVMoE - - - - 4.90% 15s 1.28% 11s 3.14% 13s 3.85% 15s
MTPOMO - - - - 5.31% 8s 1.67% 5s 3.46% 6s 4.41% 7s
ReLD-MTL - - - - 4.56% 10s 0.90% 6s 2.32% 7s 3.10% 9s
GOAL-MTL 1.20% 38s - - - 4.66% 42s - - -
URS-STL 0.11% 4s -0.09% 5s 3.22% 4s 2.01% 1.4m 4.50% 8s 1.59% 6s 2.25% 6s 3.22% 8s
URS 0.45% 4s 1.06% 5s 4.98% 4s 3.06% 1.4m 6.13% 8s 1.46% 6s 3.24% 6s 5.07% 8s
Table 3: Experimental results on seen ATSP, TSP, and CVRP.
Method ATSP100 TSP100 CVRP100
Gap Time Gap Time Gap Time
Oracle 0.00% 2m 0.00% 6m 0.00% 9.1m
Sym-NCO - 0.14% 5s 1.46% 6s
BQ 1.27% 1m 0.35% 7s 3.24% 8s
LEHD - 0.59% 2s 3.68% 2s
POMO - 0.13% 5s 1.25% 6s
ICAM 3.05% 1m 0.15% 4s 2.04% 4s
MatNet 0.95% 6.5m - -
MVMoE - - 1.65% 12s
MTPOMO - - 1.85% 6s
ReLD-MTL - - 1.42% 8s
GOAL-MTL 1.77% 1m - 4.22% 48s
URS-STL 0.62% 1.1m 0.08% 6s 1.63% 6s
URS 2.26% 1.1m 0.57% 6s 1.81% 6s
4 Experiments
In this section, we present a comprehensive evaluation of URS against both classical and neural solvers. For URS, we focus on four key aspects: (1) performance on 11 seen problems; (2) zero-shot generalization on 99 unseen problems; (3) scalability to large-scale VRP variants; and (4) results on benchmark datasets. All experiments are conducted using a single NVIDIA GeForce RTX 4090 GPU (24GB of memory) for neural solvers and an Intel Xeon Gold 6240R CPU (2.40 GHz) for classical solvers.
Table 4: Zero-shot generalization on 1K test instances of $32$ unseen VRPs ($N=100$). Due to page limit, complete experimental results (including all 99 unseen variants) are deferred to Appendix I.
Method CVRPBP CVRPBPL MDOCVRPBP MDOCVRPBPTW MDOCVRPBPL MDOCVRPBPLTW MDCVRP MDOCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 13.95% 14s 13.38% 14s 56.60% 33s 63.77% 38s 56.26% 34s 62.47% 40s 33.06% 31s 29.08% 31s
MTPOMO 13.71% 7s 13.44% 8s 42.94% 24s 47.09% 29s 41.98% 27s 46.26% 32s 23.38% 23s 27.46% 23s
ReLD-MTL 13.57% 9s 12.96% 10s 37.74% 27s 45.16% 34s 37.45% 29s 44.74% 36s 18.39% 25s 22.32% 25s
URS 12.95% 7s 12.65% 8s 24.44% 22s 26.31% 26s 24.22% 24s 26.20% 28s 15.05% 23s 22.01% 23s
Method MDCVRPL MDOCVRPTW MDOCVRPB MDOCVRPBL MDOCVRPLTW SPCTSP PDCVRP OPDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 1.6h 0.00% 1.6h 0.00% 1.6h
MVMoE 33.42% 34s 51.57% 40s 28.21% 26s 28.70% 28s 51.09% 43s - - -
MTPOMO 24.32% 25s 35.71% 29s 25.06% 20s 25.01% 21s 35.39% 31s - - -
ReLD-MTL 18.48% 28s 33.55% 33s 18.50% 21s 18.50% 22s 33.26% 35s - - -
URS 15.32% 26s 26.05% 25s 15.17% 18s 14.90% 20s 25.59% 27s -2.37% 5s -1.47% 5s 4.93% 5s
Method ACVRPL ACVRPBL ACVRPB ACVRPBTW APDTSP AMDCVRPL AMDCVRP APDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS -2.24% 1.9m 6.18% 2.4m 7.10% 2.1m 9.95% 2.2m 6.21% 1m 7.15% 6.1m 8.11% 5.3m 7.03% 1.2m
Method ACVRPLTW ACVRPBLTW ACVRPBPLTW AOCVRPBLTW ACVRPTW ACVRPBPTW AOPDCVRP AOP
Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 12.22% 2.7m 10.18% 2.4m 13.82% 2.8m 14.32% 2.5m 11.28% 2.5m 14.52% 2.6m 11.13% 1.2m -5.47% 1.2m
Table 5: Generalization on different large-scale VRP variants. All models are trained on instances of size $N=100$ and employ instance augmentation $\times$ 8. PyVRP and OR-Tools can process all instances of a given size concurrently to ensure a fair comparison.
Problem Method N = 1000 N = 2000 N = 3000 N = 4000 N = 5000
Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time
CVRPTW PyVRP 159.35(0.00%) 20m 337.15(0.00%) 40m 516.97(0.00%) 1h 618.58(0.00%) 2.7h 787.25(0.00%) 3.3h
OR-Tools 178.13(11.79%) 20m 357.18(5.94%) 40m 533.97(3.29%) 1h 647.92(4.74%) 2.7h 807.128(2.52%) 3.3h
MTPOMO 219.22(37.57%) 1.7m 488.40(44.86%) 14.6m 773.00(49.53%) 49.5m 975.61(57.72%) 1.9h OOM
MVMoE 233.53(46.55%) 1.7m 624.73(85.30%) 14.6m 1094.25(111.67%) 51.9m 1486.34(140.28%) 2h OOM
ReLD-MTL 183.31(15.04%) 1.6m 396.28(17.54%) 13.4m 613.58(18.69%) 45.7m 748.65(21.03%) 1.8h OOM
URS 172.58(8.30%) 1.3m 365.27(8.34%) 11.1m 559.03(8.14%) 30.1m 673.63(8.90%) 1.2h 853.79(8.45%) 2.3h
CVRPB PyVRP 36.52(0.00%) 20m 51.80(0.00%) 40m 75.12(0.00%) 1h 93.52(0.00%) 2.7h 112.21(0.00%) 3.3h
OR-Tools 44.23(21.11%) 20m 63.33(22.26%) 40m 88.6(17.94%) 1h 108.82(16.36%) 2.7h 128.11(14.17%) 3.3h
MTPOMO 45.92(25.77%) 1.1m 83.50(61.20%) 8.8m 136.52(81.73%) 31.4m 197.99(111.70%) 1.2h OOM
MVMoE 98.60(170.03%) 1m 250.47(383.57%) 8.1m 332.93(343.19%) 28.9m 423.19(352.49%) 1.1h OOM
ReLD-MTL 38.52(5.48%) 1m 59.21(14.30%) 8m 88.36(17.62%) 28.3m 114.88(22.83%) 1.1h OOM
URS 34.86(-4.54%) 40s 50.39(-2.72%) 5.1m 72.01(-4.15%) 17.3m 91.74(-1.91%) 41.2m 111.93(-0.24%) 1.3h
OCVRPTW PyVRP 90.91(0.00%) 20m 164.00(0.00%) 40m 224.16(0.00%) 1h 299.45(0.00%) 2.7h 367.86(0.00%) 3.3h
OR-Tools 105.98(16.58%) 20m 207.041(26.24%) 40m 280.88(25.30%) 1h 370.85(23.84%) 2.7h 447.02(21.52%) 3.3h
MTPOMO 147.14(61.85%) 1.6m 297.53(81.43%) 12.8m 437.73(95.28%) 44.3m 601.63(100.91%) 1.7h OOM
MVMoE 147.35(62.09%) 1.6m 389.60(137.57%) 12.9m 634.11(182.88%) 44.7m 899.18(200.28%) 1.7h OOM
ReLD-MTL 110.00(21.01%) 1.5m 208.60(27.20%) 12.3m 292.70(30.57%) 42m 394.42(31.72%) 1.6h OOM
URS 98.76(8.63%) 1m 178.07(8.58%) 7.6m 243.88(8.80%) 25.4m 325.64(8.74%) 1h 398.72(8.39%) 1.9h
4.1 Experimental Setup
Problem Setting
We train URS on $11$ mixed variants to ensure that most features in UDR have been seen at least twice, which prevents URS from overfitting a specific feature to a single problem. Training problems and related data generations are as follows: (1) ATSP (Kwon et al., 2021);(2) TSP (Kool et al., 2019); (3) CVRP (Kool et al., 2019); (4) ACVRP (Kwon et al., 2021; Kool et al., 2019); (5) Orienteering Problem (OP) (Kool et al., 2019); (6) PCTSP (Kool et al., 2019); (7) PDTSP (Li et al., 2021); (8) CVRPTW (Zhou et al., 2024); (9) OCVRP (Zhou et al., 2024); (10) CVRPB (Zhou et al., 2024), and (11) OCVRPTW (Zhou et al., 2024). We provide further discussion of problem selection in Appendix J.1. Each of 110 variants comprises 1,000 instances. The related oracle solvers and problem representations are detailed in Appendix D.
Model & Training Setting
We use an embedding dimension $d$ of $128$ and a feed-forward layer dimension of $512$. We set the number of attention layers in the encoder to $12$. The clipping parameter $\zeta$ is set to $50$ in Equation 10. URS is trained by the REINFORCE (Williams, 1992) algorithm, and the data is generated on the fly (see Appendix E.4 for more details). We train URS for $500$ epochs with $2\text{,}000$ batches per epoch, and each batch has a size of $128$. At each training step, we randomly select one problem from $11$ problems. We employ the AdamW optimizer (Loshchilov and Hutter, 2017) with an initial learning rate of $10^{-4}$, which is decayed by a factor of 0.1 at the 451451st epoch. Weight decay is set to $10^{-6}$. More details are provided in Appendix H.
Baseline
We compare URS with the following methods: (1)Classical Solvers: PyVRP (Wouda et al., 2024), LKH3 (Helsgaun, 2017), and OR-Tools (Perron and Furnon, 2023). We run LKH3 (Helsgaun, 2017) with $10000$ trials and 1 run (Kool et al., 2019). For PyVRP and OR-Tools, we run them on a single CPU core with time limits of 20s (Li et al., 2025a); (2)Neural Solvers: (i) specialist solvers: Sym-NCO (Kim et al., 2022), BQ (Drakulic et al., 2023), LEHD (Luo et al., 2023), ICAM (Zhou et al., 2026), POMO (Kwon et al., 2020), ELG (Gao et al., 2024), MatNet (Kwon et al., 2021), and Heter-AM (Li et al., 2021); (ii) generalist neural solvers: MTPOMO (Liu et al., 2024), MVMoE (Zhou et al., 2024), ReLD-MTL (Huang et al., 2025), and GOAL-MTL (Drakulic et al., 2025). Given the disparities in problem formulations and training settings, detailed comparisons against RouteFinder (Berto et al., 2025) and CaDA (Li et al., 2025a) are deferred to Appendix J.2.
Metrics and Inference
We report the optimality gap (Gap) and total inference time (Time) for each method. The optimality gap quantifies the discrepancy between the corresponding solvers and oracle solvers. For neural solvers, we execute the source code provided by the authors using default settings. We compare the performance of URS with relevant baselines for each problem, as well as its single-task version (i.e., URS-STL). The URS-STL variant uses exactly the same architecture and hyperparameters (see Appendix H.2 for more details). For URS, we report the best result with $\times 8$ (Kwon et al., 2020) and $\times 128$ (Kwon et al., 2021) instance augmentation for symmetric and asymmetric instances, respectively. The mark (-) indicates that the method does not support the corresponding setting.
4.2 Performance Evaluation
Performance on Seen VRPs
As presented in Table 2 and Table 3, among comparable neural solvers, URS-STL achieves the lowest gap on 8 of the 11 VRPs while offering fast inference, highlighting the strength of the URS architecture when fully specialized. Although URS-MTL exhibits a performance degradation relative to STL variants, it remains competitive across all 11 problems. Using ATSP as an example, we attribute this drop to geometric structural inequality, which likely biases the model’s learning capacity toward symmetric VRPs. While the lack of explicit constraint states yields weaker performance compared to ReLD-MTL and MVMoE, URS shows clear advantages when generalizing to unseen VRP variants (Table 4) and scaling to large-size instances (Table 5).
Generalization to Unseen VRPs
We evaluate our URS on zero-shot generalization across 99 unseen VRP variants. As presented in Table 4 and Appendix I, on 44 widely-studied VRPs with diverse constraint combinations (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions. Notably, URS merely selects the next node from a given set of feasible candidates without any additional domain knowledge, making its performance especially meaningful. Benefiting from our UDR, URS can directly address 55 less-explored VRP variants that compound asymmetry and PD relations. These results highlight URS’s strong capacity for cross-problem zero-shot generalization without any fine-tuning.
Scalability to Larger-scale VRP Variants
[…]
Results on Benchmark Dataset
We further evaluate URS on benchmark instances from CVRPLib Set-X (Uchoa et al., 2017) and CVRPLib Set-AGS (also known as Set-XXL) (Arnold et al., 2019). We compare URS against several representative specialist and generalist solvers. As presented in Table 6, URS can significantly outperform existing generalist neural routing solvers in large-scale benchmark instances. Notably, as a generalist solver, URS also greatly surpasses many representative specialist models (e.g., LEHD, BQ, and ELG). In addition, the detailed results (Appendix J.3) on the Set-X dataset show that URS still maintains its position as the best overall performance. These results further underscore the practical applicability of URS in real-world large-scale scenarios.
Table 6: Comparison on Set-AGS with $N\in[3000,7000]$. All models are trained on instances of size $N=100$.
Type Method Leuven1 Leuven2 Antwerp1 Antwerp2 Avg.gap
(N=3000) (N=4000) (N=6000) (N=7000)
Specialist LEHD greedy 16.60% 34.86% 14.66% 22.77% 22.22%
BQ greedy 18.53% 30.70% 16.48% 27.67% 23.34%
[…]
GOAL greedy OOM OOM OOM OOM -
[…]
5 Ablation Study
In this section, we conduct a comprehensive ablation study and analysis to demonstrate the effectiveness and robustness of URS. The detailed results and discussion are presented below. Note that, unless otherwise stated, the instance augmentation technique is used for the ablation studies.
Effects of Key Components
We conduct an ablation study across 11 seen VRP variants to validate the effect of key components of URS, mainly including: (1) without node-type indicator $\bm{\xi}$ of UDR; (2) without problem state $C_{t}$; (3) effects of different priors in MBM ; (4) effects of alternative decoder architectures. As detailed in Figure 4, URS maintains its competitive performance without $C_{t}$; its effectiveness is further improved by the inclusion of a node-type indicator $\bm{\xi}$. For MBM, the results reveal that three priors $\{\bm{D},\bm{D}^{\mathrm{T}},\bm{R}\}$ are complementary, and their combined use allows MBM to effectively learn the geometric and relational biases inherent in various problems, resulting in superior overall performance. Furthermore, replacing the decoder with POMO or ReLD confirms the strong cross-problem robustness of URS, while integrating the decoder introduced in Zhou et al. (2026) further improves cross-problem performance. More detailed results and analyses can be found in Appendix K.
Figure 4: Effects of different components of URS.
Table 7: Results comparing performance with and without MBM, $\mathrm{WEIGHT}(\bm{\lambda})$, and $\mathrm{BIAS}(\bm{\lambda})$ on seen and unseen variants.
MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ Avg.gap Best
Seen (11) Unseen (11) Solution
$\times$ $\checkmark$ $\checkmark$ 6.33% 10.41% 3/22
$\checkmark$ $\times$ $\times$ 4.20% 8.12% 0/22
$\checkmark$ $\times$ $\checkmark$ 3.84% 8.02% 1/22
$\checkmark$ $\checkmark$ $\times$ 3.63% 7.81% 1/22
$\checkmark$ $\checkmark$ $\checkmark$ 2.74% 7.05% 17/22
MBM & Parameter Generator
To further evaluate the effects of MBM, $\mathrm{WEIGHT}(\bm{\lambda})$, and $\mathrm{BIAS}(\bm{\lambda})$ for cross-problem zero-shot generalization, we conduct tests on 11 seen problems and 11 unseen CVRP variants widely studied in (Liu et al., 2024). As shown in Table 7, the full URS model achieves the lowest average gap on both seen and unseen problems, and provides the best solution on the majority of tasks (17/22). This shows that their synergy is essential for robust cross-problem zero-shot generalization. We provide detailed results in Appendix K.5.
Table 8: Performance comparison of ablating each UDR component (i.e., $\bm{\rho}$, $\bm{\omega}$, and $\bm{\xi}$) on seen and unseen variants. Since removing positional identifiers $\bm{\rho}$ prevents the use of instance augmentation, this augmentation technique is disabled across all results to maintain a consistent inference strategy.
$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ Avg.gap Best
Seen (11) Unseen (11) Solution
$\times$ $\checkmark$ $\checkmark$ 4.65% 18.12% 3/22
$\checkmark$ $\times$ $\checkmark$ 9.81% 18.60% 0/22
$\checkmark$ $\checkmark$ $\times$ 5.43% 18.60% 4/22
$\checkmark$ $\checkmark$ $\checkmark$ 4.17% 8.51% 15/22
Effects of Each UDR Component
To clarify the independent contribution of each UDR component, we conduct ablations by removing each component in turn. As shown in Table 8, the absence of any single component leads to notable performance degradation. A complete UDR achieves the lowest average gap, affirming the necessity of our UDR design for robust zero-shot generalization. The detailed results for each VRP variant can be found in Appendix K.6.
Hypernetwork vs. Adapter
We train URS-Adapter, a URS variant that adopts a distinct decoder for each of the 11 training problems, and compare it with the original URS with a hypernetwork on the same 11 seen problems and large-scale CVRP variants (see Appendix K.7 for detailed results). URS consistently outperforms URS-Adapter across the 11 problems seen, achieving higher parameter efficiency. This performance advantage holds consistently across large-scale CVRP variants, further solidifying our conclusion.
6 Conclusion, Limitation, and Future Work
In this work, we propose a novel RL-based URS for cross-problem zero-shot generalization. The core of URS is the proposed UDR, which broadens problem coverage by substituting problem enumeration with data unification. Its generalization capability is further enhanced by an MBM and a problem-conditioned parameter generator. Extensive experiments show that URS can produce high-quality solutions for 110 VRP variants without any fine-tuning, including 99 unseen variants, significantly broadening the problem coverage and reducing reliance on domain expertise.
Limitation and Future Work
While URS achieves impressive results across a wide range of VRP variants, its current training scheme does not fully account for the inherent disparities in problem complexity. In the future, we aim to explore more effective strategies, such as hierarchical training and difficulty-adaptive sampling, to improve training efficiency. Additionally, extending our method to accommodate more complex variants(e.g., dynamic VRP) is a promising direction for future research.
Acknowledgements
This work was supported by the National Natural Science Foundation of China (Grant No. 62476118), the Guangdong Provincial Key Laboratory of Fully Actuated System Control Theory and Technology (Grant No. 2024B1212010002), the Natural Science Foundation of Guangdong Province (Grant No. 2024A1515011759), and the Center for Computational Science and Engineering at Southern University of Science and Technology. We would like to thank the anonymous reviewers and (S)ACs of ICML 2026 for their constructive comments and dedicated service to the community. Additionally, we would like to thank Mr. Rongsheng Chen for his valuable advice and assistance during this research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Appendix A Related Work
A.1 Neural Routing Solvers with Single-Task Learning
The NCO methods have emerged as a promising paradigm for tackling various routing problems, achieving impressive empirical performance. Early seminal work based on Pointer Networks ignited interest in learning end-to-end construction policies for TSP and CVRP (Vinyals et al., 2015; Bello et al., 2016; Nazari et al., 2018). Subsequently, the Transformer-style heavy-encoder and light-decoder auto-regressive architecture has become the dominant paradigm (Kool et al., 2019), while POMO (Kwon et al., 2020) exploits permutation symmetry to enhance exploration diversity. Many advancements within this paradigm have pushed promising results on small instances (e.g., 100-node TSPs) (Xin et al., 2020; Kim et al., 2022; Xiao et al., 2024a, b; Sun et al., 2024). However, models trained on a fixed instance scale often degrade sharply when evaluated out of scale. To mitigate the poor generalization performance, recent studies introduce auxiliary local policies (Gao et al., 2024), adaptive perception across instance scales (Zhou et al., 2026), or different search space reduction methods (Fang et al., 2024; Zhou et al., 2025; Chen et al., 2025). In parallel, ”heavy decoder” designs that dynamically recompute embeddings at each construction step (Drakulic et al., 2023; Luo et al., 2023) demonstrate strong generalization on large-scale instances. To improve the performance on large-scale routing instances, complementary directions are reducing solving difficulty via problem decomposition (Zheng et al., 2024b; Ye et al., 2024; Li et al., 2025b) or adopting non-autoregressive generation augmented with additional searches (e.g., Monte Carlo tree search (MCTS) and 2-opt) (Fu et al., 2021; Sun and Yang, 2023; Qiu et al., 2022; Li et al., 2023; Ye et al., 2023). Moreover, heterogeneous vehicle routing (Li et al., 2022), asymmetric distance settings (Kwon et al., 2021), multiple VRPs (Zheng et al., 2024a), and pickup-delivery problems (Li et al., 2021; Kong et al., 2024) have each motivated tailored neural designs, which have yielded competitive results. Despite these advancements, they still train a specialist model per problem type, limiting cross-problem transfer. We instead focus on a unified neural routing solver that aims to generalize across multiple routing problems without retraining from scratch for each specification.
A.2 Neural Routing Solvers with Multi-Task Learning
To address the challenge of cross-problem generalization, growing attention has shifted to multi-task learning capable of adapting to diverse routing problems. Existing efforts can broadly fall into two categories: (1) constraint combination-based multi-task learning and (2) adapter-based fine-tuning frameworks. The former line (e.g., MTPOMO (Liu et al., 2024)) treats VRP variants as combinations of a predefined set of attributes and trains a shared backbone, achieving promising results on up to $16$ VRP variants. Follow-up work extends this paradigm by introducing mixture-of-experts (MoE) modules (Zhou et al., 2024), heavy decoder designs plus knowledge distillation (Zheng et al., 2025), and constraint-aware dual-attention (Li et al., 2025a) or updated Transformer structure (Berto et al., 2025) to improve model performance. While this enables knowledge sharing across problems, its coverage is inherently bounded by the hand-specified attribute set. The second family trains a shared backbone and performs adaptation to each variant via lightweight adapters. For instance, Lin et al. (2024) pretrain a TSP backbone and adapt to new variants through adapter-based fine-tuning. Drakulic et al. (2025) integrate edge features and adopt supervised multi-task pretraining for obtaining a powerful backbone before fine-tuning on novel tasks. Wang et al. (2025a) propose a multi-armed bandit strategy to realize more efficient multi-task training. Although these reduce retraining cost, they fail to achieve zero-shot generalization. In contrast, we focus on a constructive neural routing solver supported with zero-shot capability: once trained, it can directly generate high-quality solutions for previously unseen problems without any fine-tuning.
Appendix B Setups of VRP Variants
In this paper, we consider 110 VRP variants which may simultaneously have one or more constraints from the following categories: (1) Capacity (C); (2) Open Route (O); (3) Backhaul (B); (4) Backhaul and Priority (BP); (5) Duration Limit (L); (6) Time Windows (TW); (7) Multi-Depot (MD); (8) Prize Collecting (PC); (9) Asymmetric (A); (10) Pickup and Delivery (PD). The node coordinates are randomly sampled from the unit square when available (in symmetric geometric cases, i.e., $d_{ij}=d_{ji}$), following AM (Kool et al., 2019). Below, we provide a comprehensive description of the data generation process for each constraint.
Capacity (C)
The selection of the vehicle capacity $\mathcal{C}$ is critical, as prior work (Luo et al., 2025c) demonstrates its significant impact on the solution structure. For PDVRP variants (e.g., PDCVRP, OPDCVRP, APDCVRP, and AOPDCVRP), an excessively large $\mathcal{C}$ causes the solution to approximate that of a PDTSP instance. Conversely, an overly restrictive $\mathcal{C}$ promotes a simplistic, repetitive Pickup-Delivery sequence. To ensure structurally sound solutions, we set $\mathcal{C}$=20 for PDCVRP variants. For all other VRP variants, we adopt the convention from AM (Kool et al., 2019), setting $\mathcal{C}$=50.
Open Route (O)
The open route setting implies that vehicles are not required to return to the depot after completing their service tours. This does not alter the data generation, and it specifically affects the final cost evaluation by excluding the travel distance from the last customer back to the depot.
Backhaul (B)
The standard CVRP involves vehicles delivering goods from a depot to a set of customers with positive demands. The introduction of backhauls extends this formulation to the CVRP with Backhauls (CVRPB), where vehicles must also load goods from backhaul customers and return them to the depot, in addition to serving the standard linehaul (delivery) customers. Following the experimental setup of MTPOMO (Liu et al., 2024), we generate customer demands by sampling from a discrete uniform distribution over the integers $\{1,\ldots,9\}$. Subsequently, 20% of these nodes are randomly designated as backhaul customers, and their corresponding demand values are negative, while the remainder are classified as linehaul customers. Notably, no precedence constraints are imposed between the servicing of linehaul and backhaul customers. To ensure the feasibility of the solutions constructed by our neural solver, we adhere to the MVMoE (Zhou et al., 2024) framework’s policy that the first customer visited must be a linehaul node.
Backhaul and Priority (BP)
This constraint imposes a strict precedence requirement on the standard CVRPB, dictating that all linehaul customers must be serviced prior to any backhaul customers. Consequently, once a vehicle serves a backhaul customer, it is prohibited from visiting any subsequent linehaul nodes within the same route. If the first customer served is a backhaul node, the vehicle’s load is initialized to zero. This ensures the neural solver can construct a feasible solution.
Duration Limit (L)
Following MTPOMO (Liu et al., 2024), we impose a duration limit (i.e., the maximum length of each vehicle route) of 3 in symmetric cases. Since the node coordinates are randomly sampled from the unit square, the maximum possible distance between any two nodes is $\sqrt{2}$. The setting is sufficient to ensure a vehicle can deliver goods to at least one customer and return to the depot, thus guaranteeing the feasibility of the solution. For the setting of asymmetric cases, we provide details in the description of the following asymmetric instances.
Time Windows (TW)
Consistent with the MVMoE (Zhou et al., 2024), we set time windows of the depot to $e_{0},l_{0}]$=[0,3] and its service time to 0. The service time $s_{i}$ for each customer node is set to 0.2. For the time windows for each customer node, we generate them following the methodology outlined in MVMoE (Zhou et al., 2024).
Multi-Depot (MD)
The generation of coordinates for the multiple depots aligns with that of other nodes, with the number of depots set to 3 (Berto et al., 2025). A vehicle is required to return to the same depot it departed from. To guarantee that each starting node is considered in this multi-depot variant, we set the number of trajectories for an instance to be the product of the number of depots and the number of customer nodes, maximizing the exploration of high-quality solutions.
Prize Collecting (PC)
Under the Prize Collecting condition, each node (excluding the depot) is assigned both a prize and a penalty. A prize is collected for each visited node, while a penalty is incurred for each unvisited node. The objective is to minimize the total travel distance plus the sum of penalties for all unvisited nodes. A key constraint is that a minimum amount of total prize must be collected before the vehicle is permitted to return to the depot and terminate its route. In this context, the minimum prize requirement for the four examined problems (PCTSP, SPCTSP, APCTSP, and ASPCTSP) is standardized to 1.0. Consistent with the AM (Kool et al., 2019) framework, the prizes and penalties are generated as follows: $prize\sim\text{Uniform}\left(0,\frac{4}{n}\right)$, and $penalty\sim\text{Uniform}\left(0,3\cdot\frac{k_{n}}{n}\right)$, where $n=100$ and $k_{n}=4$ in our experiments. For the OP, the objective is to maximize collected prizes without exceeding a predefined distance limit (set to 4.0 for OP and 1.0 for AOP), where unvisited nodes carry no penalties. The prize for each node is assigned proportional to its Euclidean distance from the depot. By normalizing distances to the instance’s maximum distance and discretizing them into $[0.01,1.00]$, nodes further from the depot systematically yield higher rewards, with the depot’s prize set to zero.
Asymmetric (A)
We adopt the data generation method from MatNet (Kwon et al., 2021) to create an asymmetric distance matrix. Due to the small pairwise distances in the generated distance matrix, the parameters for duration limit and time windows set for symmetric cases are not applicable to the asymmetric problems. To ensure the constraints remain valid, we adjust the duration limit and depot end time to 0.6 and 1.0, respectively.
Pickup and Delivery (PD)
For PDP variants, we adhere to the methodology in  Li et al. (2021). For a problem of size $N$, node 0 is set to the depot, nodes $1$ to $N/2$ are designated as pickup nodes, while nodes from $N/2+1$ to $N$ are designated as delivery nodes. For each pickup node $v_{i}$, its corresponding delivery node is $v_{i+N/2}$. A strict precedence constraint is imposed in PDP cases, requiring each vehicle to visit a pickup node before its corresponding delivery node. For PDCVRP variants, the demand for each delivery node is randomly sampled from a discrete uniform distribution over $\{1,2,\ldots,9\}$. The demand for each pickup node is then set to the negative of its corresponding delivery node’s demand. To ensure the feasibility of the constraints, the vehicle capacity is set to $20$.
Appendix C Unified Data Representation
In this section, we provide a detailed description of our UDR $\bm{U}=\{\mathbf{u}_{i}\}_{i=0}^{n}$ (see Table 9). Each node $\mathbf{u}_{i}$ consists of three components, i.e., $\mathbf{u}_{i}=\{\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}\}$, where $\bm{\rho}_{i},\ \bm{\omega}_{i},\ \bm{\xi}_{i}$ are a positional identifier, unified attribute set, and node-type indicator, respectively. To distinguish linehaul/delivery from backhaul/pickup nodes, we assign negative demand values to backhaul/pickup nodes (positive for linehaul/delivery), making demand the only signed component, following  (Liu et al., 2024; Zhou et al., 2024).
Table 9: Detailed description of each component in UDR.
Attribute Name Symbol Range Description Examples
Random Identifier $\eta$ $[0,1]$ Random scalar for asymmetric problems. ATSP
Node Coordinates $\{x,y\}$ $0,1]^{2}$ 2D coordinates of the node. TSP, CVRP
Demand $\delta$ [1,1][-1,1] Node demand (++ for delivery, - for pickup). CVRP, PDCVRP
Prize $\epsilon$ $[0,1]$ Prize collected for each node. OP, PCTSP
Penalty $\mu$ $[0,1]$ Penalty incurred for each node. PCTSP
Earliest Arrival Time $e$ $[0,1]$ Earliest permissible arrival time. CVRPTW
Latest Arrival Time $l$ $[0,1]$ Latest permissible arrival time (deadline). CVRPTW
Service Time $s$ $[0,1]$ Time required for service at the node. CVRPTW
Depot - $\{0,1\}$ Indicates if the node is a depot. CVRP
Pickup Node - $\{0,1\}$ Indicates if the node is a pickup location. PDTSP
Delivery Node - $\{0,1\}$ Indicates if the node is a delivery location. CVRP, PDTSP
Node in Sub-routes - $\{0,1\}$ The solution $\pi$ can have sub-routes. CVRP
Node in Open Route - $\{0,1\}$ Vehicles need not return to the depot. OCVRP
Appendix D Multi-hot Problem Representation
Leveraging the UDR, each problem instance activates only a subset of the unified feature space. Thus, for any instance $\mathcal{G}_{i}$ from different problems, we simply mark active (non-zero) feature slots with 1 to obtain a corresponding multi-hot problem representation $\bm{\lambda}_{i}$. Different from raw feature values, it encodes presence only and provides conditional signals to the following position bias weight (see Equation 6) and adaptive decoder (see Equation 9 and Equation 10), helping the model distinguish variants and refine node selection. Here, we provide detailed representations for each problem in our experiments, and each problem is a 13-dimensional multi-hot vector.
As shown in Table 10, we train URS on eleven classic and varied routing problems to ensure that each feature in the UDR has been seen at least once, i.e., $\mathcal{P}=\{P_{i}\}_{i=1}^{11}$. We adhere to the principle of minimal redundancy and strictly align with existing work on CVRP variant selection (e.g., MVMoE (Zhou et al., 2024) and ReLD (Huang et al., 2025)): most features in UDR appear in at least two training problems, helping prevent URS from overfitting to a single problem. The only exception is the Penalty attribute, as there is structural similarity between the seen PCTSP and unseen SPCTSP. For more detailed analysis and experimental results, please see Appendix J.1.
Then, we evaluate generalization on unseen problems on different combinations of ten constraint attributes (see Appendix B), and the detailed representations of each unseen problem can be found in Table 11 (unseen well-studied VRP variants) and Table 12 (unseen less-explored VRP variants).
Table 10: The representations for seen routing problems in our experiments. The total number of problems is $1$. In training, each feature in the UDR has been seen at least once. Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
Problem Oracle RI Coord. Demand Prize Penalty EAT LAT ST Depot Pickup Delivery Sub-routes Open Route
ATSP LKH-3 $\checkmark$
TSP LKH-3 $\checkmark$
OP Compass $\checkmark$ $\checkmark$ $\checkmark$
PCTSP ILS $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PDTSP LKH-3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRP LKH-3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Frequency 2/11 9/11 6/11 2/11 1/11 2/11 2/11 2/11 9/11 2/11 7/11 6/11 2/11
Table 11: The problem representations of unseen $44$ well-studied VRP variants. Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
Problem Oracle RI Coord. Demand Prize Penalty EAT LAT ST Depot Pickup Delivery Sub-routes Open Route
CVRPL LKH-3 $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
CVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPB PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPB PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBP PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDCVRPBPTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBPL PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
MDOCVRPBPLTW PyVRP $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
SPCTSP ILS $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Table 12: The problem representations of unseen $55$ less-explored VRP variants. Here, RI represents the random identifier, and EAT, LAT, and ST indicate the earliest arrival time, the latest arrival time, and the service time, respectively.
Problem Oracle RI Coord. Demand Prize Penalty EAT LAT ST Depot Pickup Delivery Sub-routes Open Route
ACVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ACVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPB OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDCVRPBPTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBPL OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AMDOCVRPBPLTW OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
APDTSP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
APDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOPDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
PDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
OPDCVRP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
AOP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$
APCTSP OR-Tools $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
ASPCTSP N/A $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Appendix E Model Architecture and Training
E.1 Encoder
Similar to previous work (Kool et al., 2019; Kwon et al., 2020), we also use the encoder composed of $L$ stacked attention layers, where each attention layer consists of two sub-layers: an attention sub-layer and a Feed-Forward (FF) sub-layer, both of which use Instance Normalization (Ulyanov et al., 2016) and skip-connection (He et al., 2016).
The encoder $\bm{\theta}_{enc}$ with $L$ stacked attention layers transforms $H^{(0)}$ into advanced node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$. Let ${H}^{(\ell-1)}=\{\mathbf{h}_{i}^{(\ell-1)}\}_{i=0}^{n}$ denote the input to the $\ell$-th attention layer for $\ell=1,\ldots,L$. The outputs for the $i$-th node are computed as follows:
$\widetilde{\mathbf{h}}_{i}^{(\ell)}=\mathrm{IN}^{(\ell)}\left(\mathbf{h}_{i}^{(\ell-1)}+\mathrm{MBM}^{(\ell)}\left(\mathbf{h}_{i}^{(\ell-1)},H^{(\ell-1)},\bm{D},\bm{D}^{\mathrm{T}},\bm{R}\right)\right),$ (11)
$\mathbf{h}_{i}^{(\ell)}=\mathrm{IN}^{(\ell)}\left(\mathbf{\widetilde{h}}_{i}^{(\ell)}+\mathrm{FF}^{(\ell)}\left(\mathbf{\widetilde{h}}_{i}^{(\ell)}\right)\right),$ (12)
where $\mathrm{IN}(\cdot)$ denotes instance normalization (Ulyanov et al., 2016), $\mathrm{MBM}$ represents the adopted mixed bias module (see Equation 5 for details), and $\mathrm{FF}(\cdot)$ in Equation 12 corresponds to a fully connected neural network with ReLU activation. After $L$ attention layers, the final node embeddings $H^{(L)}=\{\mathbf{h}_{i}^{(L)}\}_{i=0}^{n}$ encapsulate the advanced feature representations of each node. Notably, to ensure dimensional consistency across bias channels, we normalize $\bm{D}$ and $\bm{D^{\mathrm{T}}}$ by dividing their maximum value before passing the MBM.
E.2 Decoder
In URS, all parameters $\bm{\theta}_{dec}(\bm{\lambda})$ are adaptively generated conditioned on the multi-hot problem representation $\bm{\lambda}$ (see Appendix D for detailed problem representations). $\bm{\theta}_{dec}(\bm{\lambda})$ includes two MLP networks, which are $\mathrm{WEIGHT}(\bm{\lambda})$ and $\mathrm{BIAS}(\bm{\lambda})$. $\mathrm{WEIGHT}(\bm{\lambda})$ is used to generate linear projection matrices $W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})]$. $\mathrm{BIAS}(\bm{\lambda})$ in Equation 6 is used to generate bias weights for $\mathrm{Attention}(\cdot)$(see Equation 9) and compatibility (see Equation 10).
For $\mathrm{WEIGHT}(\bm{\lambda})$, we use a multi-layer MLP hypernetwork conditioned on a 13-dimensional problem representation $\bm{\lambda}$ to generate problem-specific adaptive parameters. Note that we factor $W_{Q}(\bm{\lambda})\in\mathbb{R}^{(2d+1)\times d}$ into three vertically concatenated blocks $W_{\text{first}}(\bm{\lambda})\in\mathbb{R}^{d\times d},W_{\text{last}}(\bm{\lambda})\in\mathbb{R}^{d\times d}$ and $W_{C}(\bm{\lambda})\in\mathbb{R}^{1\times d}$, applied respectively to the first visited node $\mathbf{h}_{\pi_{1}}^{(L)}$, the last visited node $\mathbf{h}_{\pi_{t-1}}^{(L)}$, and a scalar state feature $C_{t}\in\mathbb{R}^{1}$ (see Appendix E.3 for more details). Thus $\mathrm{WEIGHT}(\bm{\lambda})$ outputs the linear projection matrices $\{W_{\text{first}}(\bm{\lambda}),W_{\text{last}}(\bm{\lambda}),W_{C}(\bm{\lambda}),W_{K}(\bm{\lambda}),W_{V}(\bm{\lambda})\}$.
Furthermore, we adopt the parameter compression technique of Ha et al. (2016) to control model size. Specifically, we first generate an advanced embedding matrix $\bm{H}_{\text{dec}}(\bm{\lambda})=[\bm{h}_{\text{dec}}^{i}(\bm{\lambda})|i=1,2,\ldots,5]\in\mathbb{R}^{5\times|\bm{\lambda}|}$ whose five row embeddings each parameterize one target projection matrix. It can be computed via three consecutive linear projections:
$\bm{H}_{\text{dec}}(\bm{\lambda})=((\bm{\lambda}W_{1}+\mathbf{b}_{1})W_{2}+\mathbf{b}_{2})W_{3}+\mathbf{b}_{3},$ (13)
where $W_{1}\in\mathbb{R}^{|\bm{\lambda}|\times d_{h}},\ W_{2}\in\mathbb{R}^{d_{h}\times d_{h}},\ W_{3}\in\mathbb{R}^{d_{h}\times(5\times|\bm{\lambda}|)},\ \mathbf{b}_{1}\in\mathbb{R}^{d_{h}},\ \mathbf{b}_{2}\in\mathbb{R}^{d_{h}},\mathbf{b}_{3}\in\mathbb{R}^{(5\times|\bm{\lambda}|)}$ are learnable parameters, $d_{h}$ is $256$ in this work, following Lin et al. (2022). Next, we use five different linear projections to map the new representation $\bm{h}_{\text{dec}}^{i}(\bm{\lambda})$ of each row to the corresponding decoder parameters. The computation can be expressed as:
$W_{\text{first}}(\bm{\lambda})=\bm{h}_{\text{dec}}^{1}(\bm{\lambda})W^{1}_{\textbf{hyper}},\ W_{\text{last}}(\bm{\lambda})=\bm{h}_{\text{dec}}^{2}(\bm{\lambda})W^{2}_{\textbf{hyper}},\ W_{C}(\bm{\lambda})=\bm{h}_{\text{dec}}^{3}(\bm{\lambda})W^{3}_{\textbf{hyper}},$ (14)
$W_{K}(\bm{\lambda})=\bm{h}_{\text{dec}}^{4}(\bm{\lambda})W^{4}_{\textbf{hyper}},\ W_{V}(\bm{\lambda})=\bm{h}_{\text{dec}}^{5}(\bm{\lambda})W^{5}_{\textbf{hyper}},$ (15)
where $W^{1}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$, $W^{2}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$, $W^{3}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(1\times d)}$, $W^{4}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$, $W^{5}_{\textbf{hyper}}\in\mathbb{R}^{|\bm{\lambda}|\times(d\times d)}$ are learnable parameters. In this way, we use a simple MLP hypernetwork $\mathrm{WEIGHT}(\bm{\lambda})$ to generate used linear projection matrices $\{W_{Q}(\bm{\lambda}),\ W_{K}(\bm{\lambda}),\ W_{V}(\bm{\lambda})\}$ conditioned on the multi-hot problem representation $\bm{\lambda}$, enabling the adaptive construction of high-quality solutions for different problems.
E.3 Problem State $C_{t}$
In this section, we detail the problems that require additional state signals. If a problem is not mentioned in this section, it means we will not apply any extra state to it, i.e., $C_{t}=0$.
For all problems with the capacity constraint, we explicitly impose only remaining load, while diverse additional constraints (e.g., time windows, duration limit, and backhaul) are enforced implicitly as $\mathcal{M}_{t}$ masks infeasible candidate nodes to guarantee valid solutions, instead of being fully enumerated in the input of the decoder. For OP and AOP, we keep track of the remaining maximum length at time $t$, and $C_{t}$ represents the remaining routing length that can be moved, and we normalize it to $[0,1]$ by dividing by the maximum length (i.e., 4.0 in OP100 and 1.0 in AOP100). The setting is the same as AM (Kool et al., 2019). For (A)PCTSP and (A)SPCTSP, $C_{t}$ is the remaining prize to collect, and we do not provide any information about the penalties, as this is irrelevant for the remaining decisions, following Kool et al. (2019). For the problem state, we provide an ablation study in Appendix K.2.
E.4 Training
Following Kwon et al. (2020), we use $n$ trajectories with distinct starting nodes in training. URS is trained by the REINFORCE (Williams, 1992) algorithm with a shared baseline (Kwon et al., 2020):
$\nabla_{\theta}\mathcal{L}(\bm{\theta})=\mathbb{E}_{p(\bm{\pi}|\mathcal{G},{\bm{\theta}})}[(f(\bm{\pi}|\mathcal{G})-\frac{1}{n}\sum_{i=1}^{n}f(\bm{\pi}^{i}|\mathcal{G}))\nabla_{\bm{\theta}}\log p_{\bm{\theta}}\left(\bm{\pi}|\mathcal{G}\right)],$ (16)
$p_{\bm{\theta}}(\bm{\pi}\mid\mathcal{G})=\prod_{t=2}^{n}p_{\bm{\theta}}(\pi_{t}\mid\mathcal{G},\pi_{1:t-1}),$ (17)
where the objective function $f(\bm{\pi}|\mathcal{G})$ represents the total reward (e.g., the negative value of tour length) of instance $\mathcal{G}$ given a specific solution $\bm{\pi}$.
Appendix F Adaptation Attention Free Module
Following Zhou et al. (2026), we implement $\mathrm{Attention}(\cdot)$ using an adaptation attention free module (AAFM) to enhance geometric pattern recognition for routing problems. Given the input $X$, the AAFM computation is then expressed as:
$Q=XW^{Q},\quad K=XW^{K},\quad V=XW^{V},$ (18)
$\mathrm{Attention}(Q,K,V,A)=\sigma(Q)\odot\frac{\exp(A)(\exp(K)\odot V)}{\exp(A)\exp(K)},$ (19)
where $W^{Q}$, $W^{K}$, and $W^{V}$ are learnable matrices. $\sigma$ denotes the sigmoid function, $\odot$ represents the element-wise product, and $A=\{a_{ij}\}$ denotes the pair-wise adaptation bias. For distance matrices, the corresponding adaptation bias $f\left(\alpha,N,d_{ij}\right)=-\alpha\cdot\log_{2}{N}\cdot d_{i,j}$. Here, $N$ denotes the total number of nodes (i.e., problem size), and $\alpha$ is generated via a lightweight network $\mathrm{BIAS}(\bm{\lambda})$ in our paper (see Equation 6). Notably, for the relation matrix, we remove the scale $N$ in the calculation of adaptation bias because the relation is scale-independent. Compared to multi-head attention (MHA) (Vaswani et al., 2017), AAFM enables the model to explicitly capture instance-specific knowledge by updating pair-wise adaptation biases while exhibiting lower computational overhead. Further details are provided in the related work mentioned above.
Appendix G Comparison between Different Attention Mechanisms
Table 13: Comparison between different attention mechanisms. Here, $N,d,m$ denote the sequence length, feature dimension, and extra hidden layer dimension of mixed-score attention, respectively. Values in parentheses represent the percentage increase or decrease relative to URS and Avg. Peak VRAM represents the average peak VRAM per instance.
Attention Mechanism Time Complexity Space Complexity #Params Avg.Peak VRAM Inference Time
Multi-Head Attention $\mathcal{O}(N^{2}d)$ $\mathcal{O}(N^{2}+Nd)$ 3.323.32M(×0.7)($\times$ 0.7) 1.871.87MB(×0.99)($\times$ 0.99) $0.92$s(×0.9)($\times$ 0.9)
Mixed-Score Attention (Parallel) $\mathcal{O}(N^{2}(d+m))$ $\mathcal{O}(N^{2}m+Nd)$ 16.1716.17M(×3.3)($\times$ 3.3) 25.6425.64MB(×13.6)($\times$ 13.6) $3.64$s(×3.7)($\times$ 3.7)
Mixed Bias Module (Ours) $\mathcal{O}(N^{2}d)$ $\mathcal{O}(N^{2}+Nd)$ 4.964.96M 1.891.89MB $0.97s$
To validate the efficiency of our proposed MBM, we have conducted a new comparative experiment between different attentions and tested them on TSP instances to ensure consistent decoding steps. Specifically, we replace our MBM with both a vanilla MHA (Vaswani et al., 2017) and the parallel mixed-score attention (MSA) used in MatNet (Kwon et al., 2021), and compare their complexities, memory overheads, and inference time. The results in Table 13 show that the increased VRAM and time overhead of MBM compared to standard MHA are small, and both maintain the same time complexity and space complexity. Compared to MatNet’s parallel MSA, our MBM drastically reduces both VRAM consumption and inference latency. These results show that MBM imposes only modest computational penalties while successfully capturing crucial problem-specific biases, which enhances cross-problem generalization while reducing architectural redundancy.
Appendix H Training and Model Settings
H.1 Multi-Task Version
Table 14: The hyperparameter settings of URS.
Hyperparameter Value
Optimizer AdamW
Clipping parameter $\zeta$ 50
Initial learning rate $10^{-4}$
Epochs of learning rate decay [451]
Factor of learning rate decay 0.1
Weight decay $10^{-6}$
The number of encoder layers $L$ $12$
Embedding dimension $d$ $128$
Feed forward dimension $512$
Dimension $d_{h}$ in Equation 13 $256$
Training scale $100$
Batches of each epoch $2,000$
Batch size $128$
Training Epochs $500$
The detailed information about the hyperparameter settings can be found in Table 14. The URS training is conducted on a single NVIDIA GeForce RTX 4090 GPU (24GB of memory). The training procedure requires only about $95$ hours ($\approx 4$ days) in total (500 epochs at approximately $12$ minutes each) to obtain a unified parameter set capable of addressing a broad class of routing problems.
H.2 Single-Task Version
URS-STL is trained on data from only a single problem type (e.g., ”URS-STL (CVRP)” is trained only on CVRP). The URS-STL variant uses exactly the same model architecture, hyperparameters (e.g., embedding dimension, number of layers, optimizer), and training setup (e.g., initial learning rate, batch size) as the multi-task URS described in Section 4.1 and Appendix H.1. The only modification we made is to the training duration. We observe that the single-task models converge significantly faster than the multi-task model. Therefore, we train URS-STL models for 300 epochs, with a learning rate decay applied at the 251st epoch. This contrasts with the multi-task URS, which is trained for 500 epochs with decay applied at the 451st epoch. This adjustment is made solely to reflect the faster convergence of the single-task training. All other hyperparameters and model architecture remain identical.
Appendix I Generalization to Unseen VRPs
To comprehensively evaluate URS’s zero-shot generalization, we test its performance on a diverse set of 99 unseen problems. This set is designed to cover both established and novel challenges, including 44 well-studied VRP variants and 55 less-explored VRP variants. The detailed results are presented in Table 15 and Table 16. On widely-studied constraints (i.e., C, O, L, TW, MD, B, BP), URS continues to deliver high-quality solutions. Notably, URS merely selects the next node from a given set of feasible candidates without any additional domain knowledge, making its performance especially meaningful. Benefiting from our UDR, URS can directly address 55 less-explored VRP variants that compound asymmetry and PD relations. Without any fine-tuning, URS still produces high-quality solutions for them, thereby strengthening its broad problem generality and deployment efficiency.
Table 15: Zero-shot generalization performance across $44$ unseen well-studied VRP variants ($N=100$).
Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 13.95% 14s 18.27% 14s 13.38% 14s 9.42% 15s 16.66% 16s
MTPOMO 13.71% 7s 17.90% 8s 13.44% 8s 9.94% 9s 17.15% 10s
ReLD-MTL 13.57% 9s 15.81% 9s 12.96% 10s 8.43% 11s 15.95% 12s
URS 12.95% 7s 16.96% 7s 12.65% 8s 10.05% 8s 18.33% 10s
Method CVRPBPTW OCVRPBPL OCVRPBPLTW CVRPL CVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 16m 0.00% 3.5h
MVMoE 16.86% 15s 17.75% 14s 8.87% 16s 0.26% 12s 1.35% 12s
MTPOMO 17.22% 9s 17.31% 8s 9.47% 10s 0.48% 6s 1.79% 6s
ReLD-MTL 16.09% 11s 15.21% 10s 8.06% 11s 0.02% 8s 1.01% 7s
URS 18.15% 9s 16.37% 8s 9.51% 9s 0.43% 7s 1.65% 6s
Method CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW MDCVRPBP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 3.5h 0.00% 6.6m
MVMoE 7.08% 15s 7.12% 12s 3.90% 15s 10.01% 15s 65.98% 33s
MTPOMO 7.41% 7s 7.34% 7s 4.37% 8s 10.50% 7s 45.77% 24s
ReLD-MTL 6.74% 8s 5.41% 7s 3.16% 9s 9.22% 8s 40.70% 27s
URS 8.94% 8s 9.47% 7s 5.12% 9s 13.72% 8s 35.44% 24s
Method MDOCVRPBP MDCVRPBPL MDOCVRPBPTW MDCVRPBPLTW MDCVRPBPTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 56.60% 33s 65.81% 36s 63.77% 38s 68.49% 42s 69.00% 39s
MTPOMO 42.94% 24s 45.98% 26s 47.09% 29s 50.45% 32s 50.74% 30s
ReLD-MTL 37.74% 27s 39.57% 29s 45.16% 34s 45.23% 37s 45.62% 35s
URS 24.44% 22s 35.07% 26s 26.31% 26s 36.49% 32s 36.71% 30s
Method MDOCVRPBPL MDOCVRPBPLTW MDCVRP MDCVRPTW MDOCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 56.26% 34s 62.47% 40s 33.06% 31s 53.64% 41s 29.08% 31s
MTPOMO 41.98% 27s 46.26% 32s 23.38% 23s 36.10% 31s 27.46% 23s
ReLD-MTL 37.45% 29s 44.74% 36s 18.39% 25s 30.90% 34s 22.32% 25s
URS 24.22% 24s 26.20% 28s 15.05% 23s 37.26% 32s 22.01% 23s
Method MDCVRPL MDCVRPB MDOCVRPTW MDOCVRPB MDCVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 33.42% 34s 21.70% 26s 51.57% 40s 28.21% 26s 20.35% 29s
MTPOMO 24.32% 25s 12.89% 19s 35.71% 29s 25.06% 20s 11.10% 20s
ReLD-MTL 18.48% 28s 9.38% 21s 33.55% 33s 18.50% 21s 8.26% 23s
URS 15.32% 26s 13.14% 19s 26.05% 25s 15.17% 18s 11.19% 21s
Method MDCVRPLTW MDOCVRPBTW MDCVRPBLTW MDOCVRPL MDCVRPBTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
MVMoE 55.55% 44s 62.47% 35s 63.18% 37s 29.33% 33s 65.90% 35s
MTPOMO 37.34% 32s 45.55% 25s 45.07% 27s 27.21% 25s 45.94% 25s
ReLD-MTL 31.56% 36s 43.36% 28s 40.57% 30s 22.05% 27s 41.28% 28s
URS 38.04% 35s 30.32% 22s 37.70% 27s 22.15% 25s 37.78% 25s
Method MDOCVRPBL MDOCVRPLTW MDOCVRPBLTW SPCTSP -
Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 1.6h -
MVMoE 28.70% 28s 51.09% 43s 62.10% 37s - -
MTPOMO 25.01% 21s 35.39% 31s 45.48% 27s - -
ReLD-MTL 18.50% 22s 33.26% 35s 43.27% 30s - -
URS 14.90% 20s 25.59% 27s 30.41% 24s -2.37% 5s -
Table 16: Zero-shot generalization performance across $55$ unseen less-explored VRP variants ($N=100$).
Method ACVRPL ACVRPBL ACVRPLTW ACVRPBLTW AOCVRPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS -2.24% 1.9m 6.18% 2.4m 12.22% 2.7m 10.18% 2.4m 37.36% 3.1m
Method AOCVRPBL AOCVRPLTW AOCVRPBLTW ACVRPBPL ACVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 17.23% 2.7m 16.35% 2.8m 14.32% 2.5m 24.03% 2.6m 13.82% 2.8m
Method AOCVRPBPL AOCVRPBPLTW ACVRPB ACVRPTW AOCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 32.50% 2.7m 16.43% 2.7m 7.10% 2.1m 11.28% 2.5m 37.63% 2.9m
Method AOCVRPTW AOCVRPB AOCVRPBTW ACVRPBTW ACVRPBP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 17.47% 2.6m 18.05% 2.5m 14.45% 2.3m 9.95% 2.2m 24.20% 2.4m
Method AOCVRPBP AOCVRPBPTW ACVRPBPTW APDTSP AMDCVRPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 32.58% 2.5m 17.45% 2.5m 14.52% 2.6m 6.21% 1m 7.15% 6.1m
Method AMDCVRPBL AMDCVRPLTW AMDCVRPBLTW AMDOCVRPL AMDOCVRPBL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 17.51% 6.5m 19.98% 8.8m 16.73% 7.5m 47.80% 8.6m 27.23% 7.1m
Method AMDOCVRPLTW AMDOCVRPBLTW AMDCVRPBPL AMDCVRPBPLTW AMDOCVRPBPL
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 24.90% 7.9m 22.84% 6.7m 35.00% 7.2m 20.50% 8.3m 38.78% 7.3m
Method AMDOCVRPBPLTW AMDCVRP AMDCVRPTW AMDOCVRP AMDCVRPB
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 24.05% 7.5m 8.11% 5.3m 20.96% 8.1m 48.02% 7.8m 17.35% 5.8m
Method AMDOCVRPTW AMDOCVRPB AMDOCVRPBTW AMDCVRPBTW AMDCVRPBP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 26.02% 7.3m 27.78% 6.5m 23.67% 6.2m 17.52% 6.8m 36.18% 6.5m
Method AMDOCVRPBP AMDOCVRPBPTW AMDCVRPBPTW PDCVRP OPDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m
URS 39.31% 6.7m 24.99% 6.9m 20.98% 7.6m -1.47% 4.3s 4.93% 4.3s
Method APDCVRP AOPDCVRP AOP APCTSP ASPCTSP
Gap Time Gap Time Gap Time Gap Time Value Time
Oracle 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m 0.00% 6.6m -
URS 7.03% 1.2m 11.13% 1.2m -5.47% 1.2m 43.79% 1.2m 2.22 1.2m
Appendix J More Experiments
J.1 Training Problem Selection
To empirically validate our selection strategy of training problems, we train two additional variants and compare them with our original URS: URS-Diff (Different Selection) and URS-Expand (Broader Problems). The first variant, URS-Diff, uses a different problem set where OCVRPTW is replaced by attribute-independent CVRPL. This modification means that the Open Route and Time Window attributes each appear in only one training problem. The second variant, URS-Expand, is trained on a broader set of 22 problems, incorporating all variants evaluated in MTPOMO (Liu et al., 2024) and MVMoE (Zhou et al., 2024).
As shown in Table 17, URS-Diff performs significantly worse on unseen problems involving Open Route and Time Window constraints. Meanwhile, URS-Expand achieves a slightly better performance than URS due to seeing more problem variants during training. However, the original URS remains highly competitive. Remarkably, despite being trained on only half the number of problem types (11 vs. 22), URS achieves the best solution on 12 out of the 22 evaluated problems.
These results demonstrate that the choice of our selected 11 problems offers a balanced trade-off, providing sufficient attribute coverage for strong generalization while maintaining training efficiency. Expanding the training set yields only marginal gains, while a poorly selected set leads to poor cross-problem generalization. In future work, we plan to explore more effective training problem selection strategies to further enhance model performance and training efficiency.
Table 17: Comparison under different training problem settings. The symbol * indicates the original problem selection, and $\Delta$ represents a different problem selection of equal number. All test problems are encompassed within URS-Expand.
Method ATSP (* $\Delta$) TSP(* $\Delta$) OP(* $\Delta$) PCTSP(* $\Delta$) PDTSP(* $\Delta$) ACVRP($\Delta$)
URS-Diff 4.52% 1.05% 0.82% 1.31% 6.09% 5.10%
URS-Expand 4.77% 0.93% 0.99% 1.41% 5.85% 5.48%
URS 2.26% 0.57% 0.45% 1.06% 4.98% 3.06%
CVRP(* $\Delta$) CVRPTW(* $\Delta$) CVRPB(* $\Delta$) OCVRP(* $\Delta$) OCVRPTW(*) CVRPL($\Delta$)
URS-Diff 1.97% 6.26% 1.80% 3.75% 15.45% 0.58%
URS-Expand 2.04% 5.62% 1.77% 3.45% 4.34% 0.65%
URS 1.81% 6.13% 1.46% 3.24% 5.07% 0.43%
OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
URS-Diff 7.55% 2.02% 2.76% 27.48% 10.25% 3.73%
URS-Expand 5.39% 1.88% 2.03% 9.13% 6.29% 3.41%
URS 9.35% 1.65% 2.67% 13.77% 9.18% 3.22%
CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW Avg.gap Best Sol.
URS-Diff 10.00% 7.56% 15.61% 27.80% 7.43% 0/22
URS-Expand 6.07% 5.47% 4.38% 9.23% 4.12% 10/22
URS 8.94% 9.47% 5.12% 13.72% 4.89% 12/22
J.2 Comparison with RouteFinder and CaDA
We strictly adhered to the problem settings established by MVMoE (Zhou et al., 2024) and ReLD (Huang et al., 2025) to ensure a fair and consistent evaluation. Notably, as shown in Table 6, URS already outperforms RF (Berto et al., 2025) and CaDA (Li et al., 2025a) on the large-scale CVRPLib benchmark dataset. In our cross-problem test, RF and CaDA are excluded due to discrepancies in the problems, including the number of trained CVRP variants (16 in their work vs. 5 in ours) and constraint settings (i.e., Backhauls and Time Windows). To ensure a fair comparison with RF and CaDA, we conduct two additional experiments:
  • We evaluate URS against CaDA and the best-performing RouteFinder variant (RF-TE) on a shared subset of both seen and unseen problems, strictly restricting the evaluation to variants with perfectly aligned constraint definitions across all methods. As shown in Table 18, when evaluated under strictly matched conditions, URS still performs best overall among comparable methods. Note that URS can provide the strongest zero-shot generalization across all shared unseen problems.
  • We retrain both RouteFinder-TE (denoted as RF-Retrain) and CaDA (denoted as CaDA-Retrain) from scratch using the exact same training settings as our URS and evaluate the three models on 48 CVRP variants. Since RF fails to train beyond 150 epochs due to the occurrence of NaN probabilities in our training setting, we evaluate all three methods at exactly 150 epochs to ensure a fair comparison. The experimental results in Table 19 show that under consistent training settings, URS still achieves the best overall performance among the three methods and provides the best solutions for the majority of problems (32/48), highlighting our strong zero-shot generalization.
Table 18: Comparison of shared seen and unseen VRP variants between URS, RF-TE, and CaDA.
Method Seen Unseen Avg.gap Best Sol.
CVRP OCVRP MDCVRP MDOCVRP MDCVRPL MDOCVRPL
RF-TE 2.06% 2.86% 63.60% 33.52% 66.75% 34.40% 33.86% 0/6
CaDA 2.17% 2.80% 40.96% 53.21% 42.40% 53.34% 32.48% 1/6
URS 1.81% 3.24% 15.04% 22.01% 15.32% 22.15% 13.26% 5/6
Table 19: Comparison between URS and the retrained RouteFinder and CaDA on all 48 CVRP variants. Note that all models are trained for 150 epochs to ensure a fair comparison.
Method CVRPTW OCVRP CVRP CVRPB OCVRPTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 8.38% 8s 6.03% 6s 3.58% 6s 4.47% 5s 8.18% 7s
CaDA-Retrain 5.28% 9s 3.22% 7s 1.92% 7s 1.61% 6s 4.42% 8s
URS 6.09% 8s 3.71% 6s 2.07% 6s 1.90% 5s 4.89% 8s
Method OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 11.61% 5s 4.73% 6s 4.77% 9s 14.17% 7s 10.11% 8s
CaDA-Retrain 8.46% 6s 4.42% 7s 4.63% 11s 10.39% 8s 10.87% 10s
URS 8.01% 6s 2.12% 6s 2.53% 9s 12.37% 7s 8.74% 8s
Method OCVRPL CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 5.98% 7s 9.74% 7s 11.58% 6s 8.34% 8s 14.46% 7s
CaDA-Retrain 5.81% 8s 7.17% 8s 12.48% 7s 8.55% 11s 15.94% 10s
URS 3.72% 7s 8.42% 7s 8.05% 6s 5.06% 8s 12.37% 8s
Method CVRPBP OCVRPBP CVRPBPL OCVRPBPTW CVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 14.40% 7s 18.31% 7s 14.01% 7s 11.43% 9s 18.70% 10s
CaDA-Retrain 13.37% 8s 18.25% 8s 16.71% 9s 9.64% 10s 20.17% 12s
URS 13.37% 7s 16.08% 7s 12.94% 8s 10.07% 9s 18.47% 10s
Method CVRPBPTW OCVRPBPL OCVRPBPLTW CVRPL MDCVRP
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 18.84% 9s 17.76% 8s 10.88% 9s 2.21% 7s 27.10% 18s
CaDA-Retrain 17.11% 10s 21.44% 9s 13.54% 12s 2.29% 8s 27.09% 23s
URS 18.63% 9s 15.42% 8s 9.57% 10s 0.66% 7s 38.17% 23s
Method MDCVRPTW MDOCVRP MDCVRPL MDCVRPB MDOCVRPTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 76.46% 25s 19.09% 18s 27.41% 34s 21.40% 15s 35.23% 22s
CaDA-Retrain 51.08% 30s 30.01% 19s 27.59% 21s 16.47% 19s 54.46% 29s
URS 38.25% 31s 21.34% 21s 39.22% 26s 20.83% 18s 28.40% 22s
Method MDOCVRPB MDCVRPBL MDCVRPLTW MDOCVRPBTW MDCVRPBLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 17.84% 15s 20.02% 29s 79.33% 40s 44.11% 19s 82.19% 23s
CaDA-Retrain 26.72% 16s 15.02% 18s 43.43% 29s 64.29% 25s 50.11% 24s
URS 15.20% 17s 17.36% 21s 39.98% 33s 25.68% 19s 33.24% 27s
Method MDOCVRPL MDCVRPBTW MDOCVRPBL MDOCVRPLTW MDOCVRPBLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 19.02% 19s 84.94% 21s 17.72% 16s 35.06% 24s 44.12% 21s
CaDA-Retrain 30.61% 20s 61.39% 25s 26.94% 17s 39.49% 29s 49.12% 25s
URS 21.56% 24s 33.58% 25s 14.84% 19s 28.60% 24s 25.84% 21s
Method MDCVRPBP MDOCVRPBP MDCVRPBPL MDOCVRPBPTW MDCVRPBPLTW
Gap Time Gap Time Gap Time Gap Time Gap Time
RF-Retrain 54.74% 21s 36.08% 21s 54.19% 23s 47.24% 26s 84.51% 30s
CaDA-Retrain 54.45% 25s 52.38% 22s 50.69% 24s 69.07% 32s 55.63% 32s
URS 41.49% 24s 24.18% 23s 41.08% 27s 24.19% 25s 33.99% 35s
Method MDCVRPBPTW MDOCVRPBPL MDOCVRPBPLTW Avg.gap Best Solutions
Gap Time Gap Time Gap Time
RF-Retrain 85.12% 28s 36.06% 23s 47.63% 28s 29.24% 3/48
CaDA-Retrain 66.78% 33s 53.48% 23s 53.06% 31s 28.45% 13/48
URS 34.55% 32s 24.10% 25s 24.32% 27s 19.01% 32/48
J.3 Detailed Results on CVRPLib Set-X Dataset
We further evaluate the generalization ability of URS on benchmark instances from CVRPLib Set-X (Uchoa et al., 2017) with scale $\in[500,1000]$. Our test results on CVRPLib Set-X datasets are presented in Table 20. All models are trained on instances of size $N=100$. In the Set-X dataset, some results for comparative models are sourced from MVMoE (Zhou et al., 2024) and RouteFinder (Berto et al., 2025). As shown in Table 20, URS maintains its position as the best overall performance across instances of varying scales. Notably, RL-based URS surpasses the SL-based representative specialist model LEHD (8.678% vs. 12.836%) and many representative multi-task neural solvers (e.g., MVMoE and RouteFinder), underscoring its practical applicability in real-world scenarios.
Table 20: Results on large-scale CVRPLib instances (Set-X) (Uchoa et al., 2017) ($500\leq N\leq 1000$). All models are trained on instances of size $N=100$.
Set-X POMO LEHD MTPOMO MVMoE/4E MVMoE/4E-L RF-MVMoE RF-TE URS
Instance Opt. Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap Obj. Gap
X-n502-k39 69226 75617 9.232% 71438 3.195% 77284 11.640% 73533 6.222% 74429 7.516% 76338 10.274% 71791 3.705% 71281 2.969%
X-n513-k21 24201 30518 26.102% 25624 5.880% 28510 17.805% 32102 32.647% 31231 29.048% 32639 34.866% 28465 17.619% 26166 8.119%
X-n524-k153 154593 201877 30.586% 280556 81.480% 192249 24.358% 186540 20.665% 182392 17.982% 170999 10.612% 174381 12.800% 175250 13.362%
X-n536-k96 94846 106073 11.837% 103785 9.425% 106514 12.302% 109581 15.536% 108543 14.441% 105847 11.599% 103272 8.884% 102969 8.564%
X-n548-k50 86700 103093 18.908% 90644 4.549% 94562 9.068% 95894 10.604% 95917 10.631% 104289 20.287% 100956 16.443% 89768 3.539%
X-n561-k42 42717 49370 15.575% 44728 4.708% 47846 12.007% 56008 31.114% 51810 21.287% 53383 24.969% 49454 15.771% 45964 7.601%
X-n573-k30 50673 83545 64.871% 53482 5.543% 60913 20.208% 59473 17.366% 57042 12.569% 61524 21.414% 55952 10.418% 54361 7.278%
X-n586-k159 190316 229887 20.792% 232867 22.358% 208893 9.761% 215668 13.321% 214577 12.748% 212151 11.473% 205575 8.018% 202645 6.478%
X-n599-k92 108451 150572 38.839% 115377 6.386% 120333 10.956% 128949 18.901% 125279 15.517% 126578 16.714% 116560 7.477% 114423 5.507%
X-n613-k62 59535 68451 14.976% 62484 4.953% 67984 14.192% 82586 38.718% 74945 25.884% 73456 23.383% 67267 12.987% 65901 10.693%
X-n627-k43 62164 84434 35.825% 67568 8.693% 73060 17.528% 70987 14.193% 70905 14.061% 70414 13.271% 67572 8.700% 66499 6.973%
X-n641-k35 63682 75573 18.672% 68249 7.172% 72643 14.071% 75329 18.289% 72655 14.090% 71975 13.023% 70831 11.226% 67005 5.218%
X-n655-k131 106780 127211 19.134% 117532 10.069% 116988 9.560% 117678 10.206% 118475 10.952% 119057 11.497% 112202 5.078% 110237 3.237%
X-n670-k130 146332 208079 42.197% 220927 50.977% 190118 29.922% 197695 35.100% 183447 25.364% 168226 14.962% 168999 15.490% 184010 25.748%
X-n685-k75 68205 79482 16.534% 72946 6.951% 80892 18.601% 97388 42.787% 89441 31.136% 82269 20.620% 77847 14.137% 75942 11.344%
X-n701-k44 81923 97843 19.433% 86327 5.376% 92075 12.392% 98469 20.197% 94924 15.870% 90189 10.090% 89932 9.776% 86038 5.023%
X-n716-k35 43373 51381 18.463% 46502 7.214% 52709 21.525% 56773 30.895% 52305 20.593% 52250 20.467% 49669 14.516% 46496 7.200%
X-n733-k159 136187 159098 16.823% 149115 9.493% 161961 18.925% 178322 30.939% 167477 22.976% 156387 14.833% 148463 9.014% 147743 8.485%
X-n749-k98 77269 87786 13.611% 83439 7.985% 90582 17.229% 100438 29.985% 94497 22.296% 92147 19.255% 85171 10.227% 83759 8.399%
X-n766-k71 114417 135464 18.395% 131487 14.919% 144041 25.891% 152352 33.155% 136255 19.086% 130505 14.061% 129935 13.563% 139371 21.810%
X-n783-k48 72386 90289 24.733% 76766 6.051% 83169 14.897% 100383 38.677% 92960 28.423% 96336 33.087% 83185 14.919% 77437 6.978%
X-n801-k40 73305 124278 69.536% 77546 5.785% 85077 16.059% 91560 24.903% 87662 19.585% 87118 18.843% 86164 17.542% 77369 5.544%
X-n819-k171 158121 193451 22.344% 178558 12.925% 177157 12.039% 183599 16.113% 185832 17.525% 179596 13.581% 174441 10.321% 171024 8.160%
X-n837-k142 193737 237884 22.787% 207709 7.212% 214207 10.566% 229526 18.473% 221286 14.220% 230362 18.904% 208528 7.635% 203457 5.017%
X-n856-k95 88965 152528 71.447% 92936 4.464% 101774 14.398% 99129 11.425% 106816 20.065% 105801 18.924% 98291 10.483% 94547 6.274%
X-n876-k59 99299 119764 20.609% 104183 4.918% 116617 17.440% 119619 20.463% 114333 15.140% 114016 14.821% 107416 8.174% 105417 6.161%
X-n895-k37 53860 70245 30.421% 58028 7.739% 65587 21.773% 79018 46.710% 64310 19.402% 69099 28.294% 64871 20.444% 58137 7.941%
X-n916-k207 329179 399372 21.324% 385208 17.021% 361719 9.885% 383681 16.557% 374016 13.621% 373600 13.494% 352998 7.236% 346556 5.279%
X-n936-k151 132715 237625 79.049% 196547 48.097% 186262 40.347% 220926 66.466% 190407 43.471% 161343 21.571% 163162 22.942% 172675 30.110%
X-n957-k87 85465 130850 53.104% 90295 5.651% 98198 14.898% 113882 33.250% 105629 23.593% 123633 44.659% 102689 20.153% 90485 5.874%
X-n979-k58 118976 147687 24.132% 127972 7.561% 138092 16.067% 146347 23.005% 139682 17.404% 131754 10.740% 129952 9.225% 125353 5.360%
X-n1001-k43 72355 100399 38.759% 76689 5.990% 87660 21.153% 114448 58.176% 94734 30.929% 88969 22.962% 85929 18.760% 77739 7.441%
Avg. Gap 29.658% 12.836% 16.796% 26.408% 19.607% 18.795% 12.303% 8.678%
Appendix K Ablation Study
In this section, we conduct a detailed ablation study and analysis to demonstrate the effectiveness and robustness of URS. Please note that, unless stated otherwise, the results presented in the ablation study reflect the best result from multiple trajectories, and we employ instance augmentation to improve performance in the ablation study. We adopt the widely used 100-node setting as our primary evaluation scenario for URS.
K.1 Effects of Node-type Indicator
Table 21: The ablation study of the node-type indicator on trained VRPs
Method ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
w/o $\bm{\xi}_{i}$ 2.14% 0.48% 0.45% 1.06% 9.28% 2.84% 3.66% 6.61% 1.57% 7.87% 5.66% 3.78%
w/ $\bm{\xi}_{i}$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
To evaluate the effectiveness of node-type indicator $\bm{\xi}$, we conduct an ablation study about including and excluding $\bm{\xi}$. Table 21 shows that adding the node-type indicator reduces the average optimality gap from 3.78% to 2.74%, with large gains on the routing problem with relatively rich instance node roles, such as PDTSP, where the coexistence of a depot and paired pickup-delivery nodes (with inherent precedence coupling) allows the explicit node-role encoding to further boost performance. Meanwhile, URS maintains a competitive performance in addressing the problem of single-node roles.
K.2 Effects of Problem State
Table 22: Performance impact of removing the problem state feature $C_{t}$
Method ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
w/o $C_{t}$ 2.09% 0.45% 0.60% 1.20% 4.59% 4.03% 2.70% 5.71% 2.64% 4.02% 4.62% 2.97%
w/ $C_{t}$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
As described in Appendix E.3, the majority of VRP variants are already solved without supplying an additional problem state $C_{t}$. To test whether the remaining use of $C_{t}$ matters, we conduct an ablation in which we eliminate it entirely, enforcing $C_{t}\equiv 0$ for all problems. The results in Table 22 show that URS maintains its competitive performance even in the absence of $C_{t}$. This indicates that explicit constraint-state features are not essential to URS effectiveness. Interestingly, on (O)CVRPTW instances, the model without any explicit constraint features outperforms the version that supplies only capacity information. We guess that capacity and time windows information are equally important, so enforcing only capacity creates a bias toward load considerations.
K.3 Effects of Mixed Bias Module
Table 23: Ablation of prior components in the MBM.
$\bm{D}$ $\bm{D^{\mathrm{T}}}$ $\bm{R}$ ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
$\checkmark$ $\times$ $\times$ 16.64% 0.44% 0.25% 0.67% 7.33% 16.62% 2.42% 6.44% 3.96% 9.44% 5.47% 6.33%
$\checkmark$ $\checkmark$ $\times$ 2.09% 0.45% 0.48% 0.82% 7.66% 9.11% 2.52% 6.89% 2.68% 5.70% 6.36% 4.07%
$\checkmark$ $\times$ $\checkmark$ 16.60% 0.46% 0.49% 0.94% 4.49% 26.22% 4.16% 6.57% 3.07% 12.23% 5.53% 7.34%
$\checkmark$ $\checkmark$ $\checkmark$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
To further assess the effectiveness of MBM, we conduct an ablation study that examines different combinations of problem-specific priors within MBM. Owing to its flexibility, the MBM allows multiple priors to be imposed on the same shared attention layer. The results in Table 23 underscore the complementary roles of the three prior components: (1) The transposed distance matrix $\bm{D^{\mathrm{T}}}$ captures directional asymmetry absent from the raw distance matrix, and its inclusion markedly stabilizes optimization on asymmetric variants; (2) Removing the pickup-delivery relation matrix $\bm{R}$ degrades performance on PDTSPs, confirming that explicit relational coupling is indispensable for these instances; (3) When all priors are fused, MBM can simultaneously learn the geometric and relational biases inherent in various problems, yielding the best overall performance.
K.4 Effects of Different Decoder Architectures
Table 24: Effect of different decoder architectures on URS cross-problem performance.
Decoder ATSP TSP OP PCTSP PDTSP ACVRP CVRP CVRPTW CVRPB OCVRP OCVRPTW Avg.gap
URS-POMO 4.83% 0.74% 0.72% 1.29% 5.45% 5.09% 1.99% 6.42% 1.83% 3.62% 5.31% 3.39%
URS-ReLD 6.48% 0.90% 0.86% 1.51% 5.81% 6.51% 2.33% 6.56% 2.33% 3.62% 5.64% 3.87%
URS 2.26% 0.57% 0.45% 1.06% 4.98% 3.06% 1.81% 6.13% 1.46% 3.24% 5.07% 2.74%
We adopt an AM (Kool et al., 2019) as our basic encoder-decoder model. Multiple decoder refinements have been proposed around AM (Kwon et al., 2020; Huang et al., 2025; Zhou et al., 2026). To assess how decoder structure impacts multi-task performance, we additionally train two URS variants whose decoders follow POMO and ReLD, denoted URS-POMO and URS-ReLD. For stronger and fairer baselines, we apply Distance-aware Attention Reshaping (DAR) (Wang et al., 2025b) to both, and all decoder parameters are generated by a hypernetwork conditioned on the problem representation $\bm{\lambda}$ (see Appendix E.2). As shown in Table 24, URS-POMO and URS-ReLD still exhibit strong cross-problem robustness, while replacing the decoder with ICAM (Zhou et al., 2026) yields a further overall improvement.
Furthermore, the results in Table 24 reveal a notable optimization challenge. By incorporating an additional Feed-Forward Network layer compared to POMO, the ReLD decoder forces the hypernetwork to map to a larger parameter space, thereby increasing its learning difficulty and degrading performance. This optimization difficulty prevents URS-ReLD from matching the performance of the original ReLD-MTL, where weights are learned directly rather than being dynamically generated. Given the outstanding performance of ReLD and other heavy-decoder paradigms in single-task optimization (Luo et al., 2023; Drakulic et al., 2023), exploring how to adapt problem-conditioned generation mechanisms for heavier decoders is a highly promising direction.
K.5 Effects of MBM & Parameter Generator on Zero-shot Generalization
Table 25: Ablation of MBM, $\mathrm{WEIGHT}(\bm{\lambda})$, and $\mathrm{BIAS}(\bm{\lambda})$ on cross-problem zero-shot generalization. The symbol (*) indicates the seen problems in training.
MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ ATSP (*) TSP (*) OP (*) PCTSP (*) PDTSP (*) ACVRP (*)
$\times$ $\checkmark$ $\checkmark$ 16.64% 0.44% 0.25% 0.67% 7.33% 16.62%
$\checkmark$ $\times$ $\times$ 4.24% 1.10% 1.04% 1.42% 6.38% 4.07%
$\checkmark$ $\times$ $\checkmark$ 3.83% 1.16% 1.09% 1.52% 6.65% 4.27%
$\checkmark$ $\checkmark$ $\times$ 2.79% 1.58% 1.39% 1.92% 7.05% 3.39%
$\checkmark$ $\checkmark$ $\checkmark$ 2.26% 0.57% 0.45% 1.06% 4.98% 3.06%
MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ CVRP (*) CVRPTW (*) CVRPB (*) OCVRP (*) OCVRPTW (*) CVRPL
$\times$ $\checkmark$ $\checkmark$ 2.42% 6.44% 3.96% 9.44% 5.47% 0.98%
$\checkmark$ $\times$ $\times$ 6.23% 7.22% 5.00% 3.99% 5.41% 2.92%
$\checkmark$ $\times$ $\checkmark$ 4.94% 6.08% 3.00% 4.61% 5.21% 4.79%
$\checkmark$ $\checkmark$ $\times$ 2.90% 6.60% 3.12% 4.47% 4.76% 1.46%
$\checkmark$ $\checkmark$ $\checkmark$ 1.81% 6.13% 1.46% 3.24% 5.07% 0.43%
MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
$\times$ $\checkmark$ $\checkmark$ 15.95% 4.18% 2.89% 18.78% 10.93% 9.51%
$\checkmark$ $\times$ $\times$ 9.54% 5.30% 2.73% 14.37% 9.80% 4.64%
$\checkmark$ $\times$ $\checkmark$ 9.58% 3.25% 3.07% 14.16% 9.91% 3.88%
$\checkmark$ $\checkmark$ $\times$ 9.92% 3.34% 3.08% 14.02% 9.92% 4.52%
$\checkmark$ $\checkmark$ $\checkmark$ 9.35% 1.65% 2.67% 13.77% 9.18% 3.22%
MBM $\mathrm{WEIGHT}(\bm{\lambda})$ $\mathrm{BIAS}(\bm{\lambda})$ CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW Avg.gap Best Sol.
$\times$ $\checkmark$ $\checkmark$ 10.74% 16.14% 5.57% 18.80% 8.37% 3/22
$\checkmark$ $\times$ $\times$ 9.67% 10.67% 5.23% 14.49% 6.16% 0/22
$\checkmark$ $\times$ $\checkmark$ 9.63% 10.42% 5.51% 13.99% 5.93% 1/22
$\checkmark$ $\checkmark$ $\times$ 9.79% 10.02% 5.81% 13.98% 5.72% 1/22
$\checkmark$ $\checkmark$ $\checkmark$ 8.94% 9.47% 5.12% 13.72% 4.89% 17/22
To empirically evaluate the necessity of these components, we have conducted tests on the 11 seen problems as well as 11 additional unseen CVRP variants widely investigated in current multi-task research (e.g., MVMoE (Zhou et al., 2024)). The results are presented in Table 25. We can observe that removing MBM leads to a significant performance drop on problems with complex geometric properties, such as ATSP, ACVRP, and PDTSP, which confirms MBM is critical for perceiving specific geometric biases. In addition, removing $\mathrm{WEIGHT}(\bm{\lambda})$ or $\mathrm{BIAS}(\bm{\lambda})$ also results in performance degradation across most tasks. In terms of overall performance, URS significantly outperforms any model where a single module is removed. The full URS model achieves the lowest average gap and provides the best solution on the majority of tasks (17/22). This demonstrates that the synergy of these components is essential for robust cross-problem generalization.
K.6 Effects of Each UDR Component
To clarify the independent contribution of each UDR component, we conduct ablations by removing each component in turn. As shown in Table 26, the absence of any single component leads to notable performance degradation. A complete UDR achieves the lowest average gap and provides the best solution on the majority of tasks (15/22), affirming the necessity of our UDR design for robust zero-shot generalization.
Table 26: The ablation study of each UDR component. The symbol (*) indicates the seen problems in training. $\bm{\rho}$, $\bm{\omega}$, and $\bm{\xi}$ represent positional identifier, unified attribute set, and node-type indicator, respectively. Since removing positional identifiers prevents the use of instance augmentation, we have excluded it from all results to ensure consistency in the inference strategy.
$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ ATSP (*) TSP (*) OP (*) PCTSP (*) PDTSP (*) ACVRP (*)
$\times$ $\checkmark$ $\checkmark$ 7.64% 1.76% 1.17% 1.96% 8.54% 7.16%
$\checkmark$ $\times$ $\checkmark$ 10.07% 1.98% 2.70% 7.05% 13.50% 13.36%
$\checkmark$ $\checkmark$ $\times$ 5.43% 1.02% 1.08% 1.86% 12.72% 5.81%
$\checkmark$ $\checkmark$ $\checkmark$ 5.51% 1.17% 0.96% 1.79% 7.71% 6.11%
$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ CVRP (*) CVRPTW (*) CVRPB (*) OCVRP (*) OCVRPTW (*) CVRPL
$\times$ $\checkmark$ $\checkmark$ 2.68% 6.88% 2.84% 4.64% 5.91% 1.29%
$\checkmark$ $\times$ $\checkmark$ 4.15% 23.71% 5.24% 7.10% 19.10% 2.70%
$\checkmark$ $\checkmark$ $\times$ 4.72% 7.67% 2.83% 9.64% 6.96% 3.33%
$\checkmark$ $\checkmark$ $\checkmark$ 2.62% 6.80% 2.69% 4.54% 6.00% 1.23%
$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ OCVRPB CVRPBL CVRPLTW OCVRPBTW CVRPBLTW OCVRPL
$\times$ $\checkmark$ $\checkmark$ 11.87% 3.12% 3.34% 34.34% 44.03% 4.63%
$\checkmark$ $\times$ $\checkmark$ 15.02% 5.47% 19.53% 30.88% 29.64% 6.98%
$\checkmark$ $\checkmark$ $\times$ 26.06% 2.69% 4.13% 34.62% 26.55% 9.68%
$\checkmark$ $\checkmark$ $\checkmark$ 11.54% 3.03% 3.38% 15.47% 10.70% 4.55%
$\bm{\rho}$ $\bm{\omega}$ $\bm{\xi}$ CVRPBTW OCVRPBL OCVRPLTW OCVRPBLTW Avg.gap Best Sol.
$\times$ $\checkmark$ $\checkmark$ 43.99% 11.99% 6.02% 34.66% 11.38% 3/22
$\checkmark$ $\times$ $\checkmark$ 29.34% 14.84% 19.22% 31.02% 14.21% 0/22
$\checkmark$ $\checkmark$ $\times$ 28.60% 25.96% 7.00% 35.93% 12.01% 4/22
$\checkmark$ $\checkmark$ $\checkmark$ 10.51% 11.68% 6.08% 15.38% 6.34% 15/22
K.7 Hypernetwork vs. Adapter
We train URS-Adapter, a URS variant that adopts a distinct decoder for each of the 11 variants, and compare it with the original URS using a hypernetwork on 11 seen problems (see Table 27) and large-scale CVRP variants (see Table 28). URS consistently outperforms URS-Adapter across 11 seen problems, achieving higher parameter efficiency. This performance advantage holds consistently on large-scale CVRP variants, which further solidifies our conclusion.
Table 27: Performance comparison between URS-Hypernet and URS-Adapter on 11 seen VRP variants.
Method OP PCTSP PDTSP ACVRP
Gap Time Gap Time Gap Time Gap Time
URS-Adapter 0.88% 4s 1.47% 4s 5.69% 4s 3.83% 1.5m
URS 0.45% 4s 1.06% 4s 4.98% 4s 3.06% 1.5m
Method CVRPTW CVRPB OCVRP OCVRPTW
Gap Time Gap Time Gap Time Gap Time
URS-Adapter 6.34% 8s 3.21% 5s 4.06% 6s 4.94% 7s
URS 6.13% 8s 1.46% 5s 3.24% 6s 5.07% 7s
Method ATSP TSP CVRP Average Best
Gap Time Gap Time Gap Time Gap Solution
URS-Adapter 2.97% 1.2m 0.82% 4s 2.85% 6s 3.37% 1/11
URS 2.26% 1.2m 0.57% 4s 1.81% 6s 2.74% 10/11
Table 28: Generalization comparison between URS and URS-Adapter across diverse large-scale CVRP variants.
Problem Method N = 1000 N = 2000 N = 3000 N = 4000 N = 5000
Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time Obj.(Gap) Time
CVRPTW PyVRP 159.35(0.00%) 20m 337.15(0.00%) 40m 516.97(0.00%) 1h 618.58(0.00%) 2.7h 787.25(0.00%) 3.3h
URS-Adapter 173.68(8.99%) 1.3m 367.22(8.92%) 10.8m 564.55(9.21%) 30.8m 677.91(9.59%) 1.2h 860.63(9.32%) 2.3h
URS 172.58(8.30%) 1.3m 365.27(8.34%) 11.1m 559.03(8.14%) 30.1m 673.63(8.90%) 1.2h 853.79(8.45%) 2.3h
CVRPB PyVRP 36.52(0.00%) 20m 51.80(0.00%) 40m 75.12(0.00%) 1h 93.52(0.00%) 2.7h 112.21(0.00%) 3.3h
URS-Adapter 42.71(16.96%) 43s 62.73(21.11%) 5.5m 85.72(14.11%) 17.5m 107.17(14.59%) 41.9m 127.93(14.01%) 1.3h
URS 34.86(-4.54%) 40s 50.39(-2.72%) 5.1m 72.01(-4.15%) 17.3m 91.74(-1.91%) 41.2m 111.93(-0.24%) 1.3h
OCVRPTW PyVRP 90.91(0.00%) 20m 164.00(0.00%) 40m 224.16(0.00%) 1h 299.45(0.00%) 2.7h 367.86(0.00%) 3.3h
URS-Adapter 105.96(16.56%) 1.1m 203.52(24.10%) 9.9m 287.70(28.35%) 28.6m 391.06(30.59%) 1h 483.19(31.35%) 1.9h
URS 98.76(8.63%) 1m 178.07(8.58%) 7.6m 243.88(8.80%) 25.4m 325.64(8.74%) 1h 398.72(8.39%) 1.9h
Appendix L Licenses for Used Resources
Table 29: List of licenses for the codes and datasets we used in this work.
Resource Type Link License
PyVRP (Wouda et al., 2024) Code https://github.com/PyVRP/PyVRP MIT License
LKH3 (Helsgaun, 2017) Code http://webhotel4.ruc.dk/~keld/research/LKH-3/ Available for academic research use
OR-Tools (Perron and Furnon, 2023) Code https://github.com/google/or-tools Apache-2.0 License
POMO (Kwon et al., 2020) Code https://github.com/yd-kwon/POMO/tree/master/NEW_py_ver MIT License
Sym-NCO (Kim et al., 2022) Code https://github.com/alstn12088/Sym-NCO Available for any non-commercial use
LEHD (Luo et al., 2023) Code https://github.com/CIAM-Group/NCO_code/tree/main/single_objective/LEHD Available for any non-commercial use
BQ (Drakulic et al., 2023) Code https://github.com/naver/bq-nco CC BY-NC-SA 4.0 license
ICAM (Zhou et al., 2026) Code https://github.com/CIAM-Group/ICAM MIT License
ELG (Gao et al., 2024) Code https://github.com/gaocrr/ELG MIT License
MatNet (Kwon et al., 2021) Code https://github.com/yd-kwon/MatNet/tree/main MIT License
Heter-AM (Li et al., 2021) Code https://github.com/jingwenli0312/Heterogeneous-Attentions-PDP-DRL MIT License
MTPOMO (Liu et al., 2024) Code https://github.com/FeiLiu36/MTNCO MIT License
MVMoE (Zhou et al., 2024) Code https://github.com/RoyalSkye/Routing-MVMoE MIT License
RouteFinder (Berto et al., 2025) Code https://github.com/ai4co/routefinder MIT License
CaDA (Li et al., 2025a) Code https://github.com/CIAM-Group/CaDA MIT License
ReLD (Huang et al., 2025) Code https://github.com/ziweileonhuang/reld-nco MIT License
GOAL (Drakulic et al., 2025) Code https://github.com/naver/goal-co Available for any non-commercial use
CVRPLib Set-X (Uchoa et al., 2017) Dataset http://vrp.galgos.inf.puc-rio.br/index.php/en/ Available for academic research use
CVRPLib Set-AGS (Arnold et al., 2019) Dataset http://vrp.galgos.inf.puc-rio.br/index.php/en/ Available for academic research use
We list the existing codes and datasets in Table 29, all of which are open-source resources for academic use.